### v1

# Learning to Control Unknown Strongly Monotone Games

Abstract
Consider $N$ players each with a $d$-dimensional action set. Each of the players’ utility functions includes their reward function and a linear term for each dimension, with coefficients that are controlled by the manager. We assume that the game is strongly monotone, so if each player runs gradient descent, the dynamics converge to a unique Nash equilibrium (NE). The NE is typically inefficient in terms of global performance. The resulting global performance of the system can be improved by imposing $K$-dimensional linear constraints on the NE. We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE. However, this requires knowing the players’ reward functions and their action sets. Obtaining this game structure information is infeasible in a large-scale network and violates the users’ privacy. To overcome this, we propose a simple algorithm that learns to shift the NE of the game to meet the linear constraints by adjusting the controlled coefficients online. Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets. We prove that our algorithm, which is based on two time-scale stochastic approximation, guarantees convergence with probability 1 to the set of NE that meet target linear constraints. We then provide a mean square convergence rate of $O(t^{-1/4})$ for our algorithm. This is the first such bound for two time-scale stochastic approximation where the slower time-scale is a fixed point iteration with a non-expansive mapping. We demonstrate how our scheme can be applied to optimizing a global quadratic cost at NE and load balancing in resource allocation games. We provide simulations of our algorithm for these scenarios.
1 Introduction
Large-scale networks such as in communication, transportation, energy, and computing consist of myriad of agents taking local decisions. Agents do not see the bigger picture therefore optimize their local objectives instead. The result is often an inefficient equilibrium. Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective. However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy. As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5, 6, 7, 8, 9].
Game theory is a natural way to tackle distributed decision-making [10, 11, 12]. The game structure formalizes the effect of the interaction between users. The players may be cooperative, or they may be selfishly interested in maximizing their reward. In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they have no information regarding the reward functions of their peers. A Nash equilibrium (NE) aims to predict the outcome of such a repeated interaction. In monotone games [13, 14, 15], which are our focus here, simple distributed algorithms such as gradient descent of players on their reward functions are guaranteed to converge to a NE.
A NE does not optimize global objectives and can lead to poor performance. By controlling parameters in the utility functions of the players, the manager can aim to change the game such that the NE is efficient. In general, setting utility parameters to optimize performance at the NE is intractable. Therefore, in this paper, we focus on the task of imposing linear constraints on the resulting NE of the game the manager controls. Linear constraint are common in many optimization formulations. When optimizing a global quadratic cost, linear constraint can describe the first-order optimality condition for the NE. In resource allocation, linear constraints can describe what should be the ideal loads at each resource at NE.
Even when setting utility parameters to impose linear constraints at the NE is tractable, it requires knowing the reward functions of players and their action sets. This would require the manager to collect this information from the players, and then solve a large-scale optimization problem. This centralized approach violates the privacy of the users and is infeasible in large-scale networks. Even worse, the manager would need to collect the parameters periodically if they are time-varying, and then solve the large-scale optimization problem again.
To converge to a NE that meets target linear constraints in an unknown strongly monotone game, we propose an online learning approach. In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback. Such feedback is easy to monitor in practice and also maintains user privacy. Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
We first prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1. We further show that our algorithm has a mean square error rate of $O(t^{-1/4})$ for the linear constraint. Due to the convex and compact action sets of the players, our slower time-scale is a fixed point iteration with a non-expansive mapping as opposed to the stronger contractive mapping assumed to analyze convergence rates in the stochastic approximation literature [16, 17]. Our bound is the first of this kind for a two time-scale stochastic approximation scheme with a non-expansive mapping. Apart from tools from stochastic approximation, we employ variational inequalities and the Krasnosel’skiĭ–Mann algorithm in our analysis. Therefore, our simple algorithm offers a theoretically principled game control scheme for strongly monotone games.
1.1 Related Work
Distributed optimization [18, 19, 20] is concerned with optimizing a target objective distributedly over a network of agents. However, these agents are non-interacting computational units so there is no game structure between them. To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward function, which is infeasible. Moreover, distributed optimization entails significant communication overhead between agents.
Distributed algorithms that converge to a NE have been studied in [21, 22, 23], including in monotone games [24, 25, 26]. The players in our game use simple stochastic gradient play to converge to a NE. However, if the game is not controlled, the NE is typically inefficient from a global point of view. Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
Compared to mechanism design approaches [27, 28, 29], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only receives the sum of actions (loads) of all players.
Our approach resembles the idea of a Stackelberg game [30, 31, 32, 33], where the manager is the leader and the players are the followers (usually a single follower is considered). However, game-theoretic algorithms converge to a NE that can be inefficient from an optimization point of view. Our work provides a mechanism that provably shifts the unique NE to a point that satisfies target linear constraints that can force the NE to be efficient (e.g., first-order optimality condition for a global quadratic objective).
Our learning framework can be viewed as a bandit learning problem where the manager is playing against a bandit that is not stochastic nor fully adversarial [34, 35]. Instead, the rewards of this bandit are generated by the game dynamics. This “game bandit” also introduces a new type of noise. The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints. Hence, the type of feedback that the manager needs is about the behavior at NE. However, the dynamics are not at NE every turn. In fact, by changing the prices the manager perturbs the system which complicates and delays the convergence to the unique NE. Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
The most closely related literature to our work is that on intervention and control of games [36, 37, 38, 39, 40, 41, 42, 43, 44]. In [45] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE. This is also the assumption in [46, 40], that considered the general approach of game-theoretic control. In [36], an aggregative game where the cost is a linear function of the aggregated action was considered. It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads. In [47, 48, 49], a similar approach to ours was considered for discrete congestion games, where the reward function of all players is the same convex function of the number of players that share the chosen resource. The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets. For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., their geometric locations). Additionally, the pricing scheme in [47, 48, 49] requires the manager to know the reward functions. Control of dynamic Markov potential games was considered in [41], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
There have been works obtaining convergence rate for contractive two time-scale stochastic approximation [50, 16]. These works assume that both the time-scales are fixed point iterations with contractive mappings. We consider a more general case where the slower timescale has a non-expansive mapping.
A preliminary conference version, limited to load balancing in resource allocation, was published in [51]. In this paper, we extend the framework in [51] to linear constraint in strongly monotone games. In particular, the NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$. Our convergence guarantees are with probability 1 instead of in L2 like in [51]. Last but not least, we provide a finite time convergence error bound. Our novel results require new analysis based on non-expansive mappings that has limited overlap with the analysis in [51].
1.2 Outline
The rest of the paper is organized as follows. In Section 2 we formulate our game control problem and detail our assumptions. In Subsection 2.1 we detail two applications that fall under our general model. In Section 3 we describe the algorithm that the manager uses to control the game and state our main results regarding its performance guarantees. In Section 4 we break the proofs of our main results into lemmas. In Section 5 we provide numerical simulations for the application described earlier. Then, In Section 6 we conclude the paper and propose future research directions.
1.3 Notation
[…]
2 Problem Formulation
Consider a set of players $\mathcal{N}=\{1,2,\ldots,N\}$, where player $n$ chooses an action $\boldsymbol{x}_{n}=(x_{n}^{1},\ldots,x_{n}^{d})\in\mathcal{X}_{n}$. Here $\mathcal{X}_{n}\subset\mathbb{R}^{d}_{+}$ is a convex and compact set such that $\boldsymbol{0}\in\mathcal{X}_{n}$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$. We use $\boldsymbol{x}\in\mathcal{X}$ to denote the concatenation of all action vectors. Then the utility for each player is defined as:
$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{d}x_{n}^{d}.$ (1)
Here $r_{n}(\boldsymbol{x})$ is the reward function for player $n$. We assume that $r_{n}(\boldsymbol{x})$ is twice continuously differentiable for each $n$ and is greater than or equal to 00 for all $\boldsymbol{x}\in\mathcal{X}$. The vector $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ denotes the ‘controlled coefficients’ controlled by the game manager. We use the same vector notation for $\boldsymbol{\beta}$ and $\boldsymbol{\beta}_{n}$.
Figure 1: Control of an unknown game
The key assumption about the game structure is that it is a strongly monotone game. For this, we define the gradient operator $F(\cdot):\mathbb{R}^{Nd}\mapsto\mathbb{R}^{Nd}$ as
$F(\boldsymbol{x})\coloneqq(\nabla_{\boldsymbol{x}_{1}}r_{1}(\boldsymbol{x}_{1},\boldsymbol{x}_{-1}),\ldots,\nabla_{\boldsymbol{x}_{N}}r_{N}(\boldsymbol{x}_{N},\boldsymbol{x}_{-N})).$ (2)
Assumption 1.
There exists $\mu>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
$\langle\boldsymbol{y}-\boldsymbol{x},F(\boldsymbol{y})-F(\boldsymbol{x})\rangle\leq-\mu\|\boldsymbol{y}-\boldsymbol{x}\|^{2}.$ (3)
The above assumption means that the game is a strongly monotone game with parameter $\mu>0$. Monotone games are a well-studied class of games [52]. The linear term in (1) does not affect the strongly monotone property of the game since the gradient vector becomes $F(\boldsymbol{x})-\boldsymbol{\beta}$. Since it is a strongly monotone game, there exists a unique pure NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ corresponding to each vector $\boldsymbol{\beta}$, defined as:
Definition 1.
A strategy profile $(\boldsymbol{x^{*}}_{n}(\boldsymbol{\beta}),\boldsymbol{x^{*}}_{-n}(\boldsymbol{\beta}))$ is a pure Nash equilibrium (NE) corresponding to coefficients $\boldsymbol{\beta}$, if $u_{n}(\boldsymbol{x}_{n}^{\boldsymbol{*}}(\boldsymbol{\beta}),\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})\geq u_{n}(\boldsymbol{x}_{n},\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})$, for all $\boldsymbol{x}_{n}\in\mathcal{X}$ and all $n\in\mathcal{N}$.
We make the following additional assumption about the game, which says that the gradient function is Lipschitz continuous. It is merely a technical assumption, commonly taken along with the strongly monotone condition [53, 54].
Assumption 2.
There exists $L>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
$\|F(\boldsymbol{y})-F(\boldsymbol{x})\|\leq L\|\boldsymbol{y}-\boldsymbol{x}\|.$ (4)
Our game evolves in discrete time, with players taking actions $\boldsymbol{x}_{n,t}$, given the controlled coefficients $\boldsymbol{\beta}_{n,t}$, at time $t$. The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the objective of the game manager (discussed below (6)). Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., the game is uncontrolled), then the players will eventually converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
In many scenarios, the players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive (i.e., bandit feedback) or other data they collect. The gradient can then be approximated from this data using regression or other supervised machine learning techniques [55]. These learning schemes make the players’ actions stochastic which adds another type of noise to our analysis. We therefore assume that agents observe noisy variants of their gradients, where the noise is a martingale difference sequence with a bounded second moment.
Assumption 3.
Each player $n$ observes $\boldsymbol{g}_{n,t-1}$ at time $t$, where $\boldsymbol{g}_{n,t-1}=\nabla_{\boldsymbol{x}_{n,t-1}}r_{n}(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{n,t}$. Then $\boldsymbol{g}_{t-1}=F(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{t}$, where $\boldsymbol{M}_{t}$ satisfies
$\mathbb{E}[\boldsymbol{M}_{t}\mid\mathcal{F}_{t-1}]=0\;\;\;\textrm{and}\;\;\;\mathbb{E}[\|\boldsymbol{M}_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq M_{c}.$ (5)
Here $\mathcal{F}_{t}$ is the filtration that summarizes the past, i.e.,
$\mathcal{F}_{t}=\sigma(\boldsymbol{x}_{\tau},\boldsymbol{\beta}_{\tau}\mid\forall\tau\leq t).$
The objective of the manager is to steer the players’ NE towards a point that satisfies a set of $K$ linear constraints, which can together be represented in vector form as
$\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}.$ (6)
Here $\boldsymbol{A}\in\mathbb{R}^{K\times Nd}$ and $\boldsymbol{\ell^{*}}\in\mathbb{R}^{K}$. We assume that, at each time $t$, the manager only observes the constraint violation, i.e., the vector $(\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}})$. These linear constraints can have different meanings for different applications. For resource allocation, the players’ actions are how much to use of each resource and the manager wants to meet the target loads. When the global objective is a quadratic cost, the constraint violation vector is just the gradient of the global quadratic cost.
2.1 Applications
In this section, we show how our framework can be applied in two different applications.
2.1.1 Quadratic Global Objective
For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global objective of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$ which they wish to minimize. All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$. Hence, using our approach, we can set $A=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a local optimum.
2.1.2 Weighted Resource Allocation Games
In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users. Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol. In the electricity grid, the manager can be the utility company running DSM to cut the production costs of energy at peak hours. In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network. Other examples are load balancing in data-centers [56] and control of parking resources [57]. For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources. We have $d$ constraints i.e., $K=d$ with the idea of each linear constraints being to meet the target loads for each resource $i$, i.e.,
[…]
The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize for application-specific objectives such as system efficiency, regulations, and operation costs. The controlled coefficients $\boldsymbol{\beta}$ can be thought of as incentives to use a resource in this application. The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
A special case of this is resource allocation without weights, studied in [51]. In this case the linear constraints are
[…]
Generalizing to a weighted sum of actions allows modeling of cases such as power control in wireless networks and load balancing in data-centers. In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains. In data-centers, different users send jobs of different sizes to the different servers (resources), so their actions (how many jobs to send where) affect the loads differently.
3 Game Control Algorithm
Algorithm 1 Online Game Control with Bandit Feedback
Initialization: Let $\boldsymbol{x}_{0}\in\mathcal{X}$ and $\boldsymbol{\alpha}_{0}\in\mathbb{R}^{K}$.
Input: Target vector $\boldsymbol{l}^{*}$.
For each turn $t\geq 1$ do
  1. 1.
    The manager observes the vector $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$.
  2. 2.
    Each player $n$ computes $\boldsymbol{\beta}_{n,t-1}=\boldsymbol{A}^{\intercal}_{n}\boldsymbol{\alpha}_{t-1}$ and updates its action using gradient ascent:
    $\boldsymbol{x}_{n,t}=\Pi_{\mathcal{X}_{n}}\left(\boldsymbol{x}_{n,t-1}+\eta_{t-1}\left(\boldsymbol{g}_{n,t-1}-\boldsymbol{\beta}_{n,t-1}\right)\right)$ (7)
    where $\Pi_{\mathcal{X}_{n}}$ is the Euclidean projection into $\mathcal{X}_{n}$.
  3. 3.
    The manager updates the controlled input using
    $\boldsymbol{\alpha}_{t}=\boldsymbol{\alpha}_{t-1}+\epsilon_{t-1}(\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}).$ (8)
  4. 4.
    The manager broadcasts $\boldsymbol{\alpha}_{t}$ to the agents.
End
Algorithm 1 details the combined scheme of the manager’s iteration and the players’ behavior (modeled or designed). We discuss the key details of the algorithm in this section.
The manager controls the controlled coefficients $\boldsymbol{\beta}$ in order to steer the dynamics towards a desirable point. But instead of directly performing computations on $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ and communicating it to agents, the manager uses the “control input” $\boldsymbol{\alpha}\in\mathbb{R}^{K}$, where $K$ is the number of linear constraints, which can be much smaller than $Nd$ depending on the application. After computing $\boldsymbol{\alpha}_{t}$ at time $t$, the manager broadcasts $\boldsymbol{\alpha}_{t}$ to the agents who then compute their $\boldsymbol{\beta}_{n,t}$ using $\boldsymbol{\beta}_{n,t}=\boldsymbol{A}^{\intercal}_{n}\boldsymbol{\alpha}$. Here $\boldsymbol{A}^{\intercal}_{n}$ denotes the $d$ rows corresponding to player $n$ in the matrix $\boldsymbol{A}^{\intercal}$. Our algorithm requires each player to know only its corresponding $d$ rows in $\boldsymbol{A}^{\intercal}$. Note that $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$. Depending on the application, this row can represent local parameters known to the player (e.g., the size of the jobs the player sends to the server), or that the manager sends these parameters to players at the beginning of the algorithm.
The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$. To overcome this, we propose an online learning approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$. We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$. This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
Our algorithm can be thought of as a two time-scale stochastic approximation algorithm. In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$. In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
This algorithm can also be thought of as an iterative approach to solving the constrained optimization problem using Lagrange multipliers. We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$. The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied at all times as that will require communication between agents, and hence the manager uses iterative approach to solve the constrained optimization problem. Similar approaches have been used in [58, 59] for different problems.
Our players run stochastic gradient descent to optimize their rewards, and therefore are myopic and do not look to manipulate the manager using dynamic strategies. This behavior can be encountered with selfish players in a large-scale network or with cooperative players.
In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [60, 61, 15]. In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps. Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
In a cooperative setting, the behavior of the players is not something to model but to design [62]. The algorithm is used as a distributed protocol programmed into the devices in the network (e.g., WiFi). The challenge is then not to control the selfish behavior of the players, but to guide them towards a more efficient NE they do not have enough information to find otherwise (i.e., players do not know which $\boldsymbol{\alpha}$ would satisfy the linear constraints at NE). For this purpose, gradient play is an efficient distributed algorithm, that can lead to an efficient NE when combined with the manager’s algorithm in (8).
The manager’s algorithm in (8) is robust against strategic manipulations of a small group of players since each player has a negligible effect on $\boldsymbol{A}\boldsymbol{x}_{t-1}$, which is its only way to impact (8). It is not clear if even a large group of players can manipulate (8) to increase their accumulated reward over time since any manipulated change to $\boldsymbol{\alpha}_{t}$ would be temporary and incurs losses to the players. Moreover, truthful reporting is not an issue in our algorithm since the players do not report anything, and the manager cannot observe their actions. Instead, the manager only needs to measure the instantaneous constraint violation
3.1 Performance Guarantees
The manager wants to converge to a suitable $\boldsymbol{\alpha}$ such that the linear constraints in (6) are satisfied at the NE corresponding to $\boldsymbol{\alpha}$, i.e., converge to the set
$\mathcal{N}_{opt}=\{\boldsymbol{\alpha}\mid\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}\},$
where $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ denotes the NE corresponding to $\boldsymbol{\alpha}$.
We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is is just Slater’s condition [63] for convex optimization.
Assumption 4.
The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is a continuously differentiable convex function and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
We also take the following standard assumptions on our stepsize sequences [16].
Assumption 5.
The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ which satisfy:
  1. (a)
    $\{\eta_{t}\},\{\epsilon_{t}\}$ are monotonically decreasing sequences that are non-summable, but square summable.
  2. (b)
    $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$.
Our first result shows that $\boldsymbol{\alpha}_{t}$ converges to $\mathcal{N}_{opt}$ and hence $\boldsymbol{A}\boldsymbol{x}_{t}$ converges to $\boldsymbol{\ell^{*}}$.
Theorem 1.
Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and
$\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}},$
with probability 1.
The following theorem gives finite-time convergence guarantees for our algorithm.
Theorem 2.
Under Assumptions 1-5,
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right).$
Therefore the best rate is obtained for $\eta=0.5$ and $\epsilon=0.75$:
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-1/4}\right).$
The above theorem holds for all $t$ if we choose $T_{1}$ and $T_{2}$ to be sufficiently large and then $\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right)$ for all $t$, for some $C>0$.
4 Convergence Analysis
In this section, we detail our proof strategy for Theorem 1 and 2. The proofs for the lemmas below and the main theorems are given in Appendix I and II respectively.
Throughout the proofs, we use the notation $F(\boldsymbol{x},\boldsymbol{\alpha})=F(\boldsymbol{x})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$, which is the gradients vector of the game given the controlled input $\boldsymbol{\alpha}$.
Our first lemma shows that the desired set $\mathcal{N}_{opt}$ is non-empty. The proof for this lemma follows from the strongly monotone nature of the game and Slater’s conditions (Assumption 4).
Lemma 3.
Under Assumptions 1 and 4, there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, i.e., the set $\mathcal{N}_{opt}$ is non-empty.
One of the main technical challenges in our analysis stems from the constrained actions represented by the convex and compact action sets. Indeed, with no constraints on the actions, and under certain assumptions on the matrix $\boldsymbol{A}$, we can show that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is strongly monotone instead of co-coercive (Lemma 4 (b)). This in turn shows that the mapping $g(\boldsymbol{\alpha})$ in Definition 2 is contractive. In this simplified case, both time-scales are fixed point iterations with contractive mappings and results from [16] can be applied. Therefore, the action sets call for a novel two time scale convergence analysis.
4.1 Non-expansive Mapping
In this subsection, we prove the non-expansiveness of the following mapping that is central to our analysis.
Definition 2.
Define the map $g(\cdot):\mathbb{R}^{K}\mapsto\mathbb{R}^{K}$ as
$g(\boldsymbol{\alpha})=\boldsymbol{\alpha}+\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}}),$
where $\gamma=\min\{\mu/\|\boldsymbol{A}\|^{2},1\}$ is a constant chosen to satisfy $02\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2}>$ 0. Note that a fixed point of $g(\boldsymbol{\alpha})$ will satisfy our objective of $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$.
To establish the non-expansiveness of $g$, we first need to show two important properties of the NE corresponding to each $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. The first property is that $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is Lipschitz continuous in $\boldsymbol{\alpha}$ and the second property is that the weighted load vector $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is co-coercive in $\boldsymbol{\alpha}$. The proofs for these two results follow from the strongly monotone nature of the game by using properties of variational inequalities (VI).
Lemma 4.
Let $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ be the unique NE given $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. Then,
  1. (a)
    There exists a constant $L_{0}$ such that for all $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$,
    $\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|\leq L_{0}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|.$ (9)
  2. (b)
    For all $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$,
    $\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\rangle$
    $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq-\mu\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2},$ (10)
    where $\mu$ is the monotonicity parameter of the game.
The next lemma shows that the map $g(\cdot)$ is non-expansive.
Lemma 5.
For any $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$, the function $g(\cdot)$ satisfies,
$\|g(\boldsymbol{\alpha}_{2})-g(\boldsymbol{\alpha}_{1})\|\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|,$ (11)
i.e., the map $g(\cdot)$ is non-expansive.
The next lemma shows that the iterate $\boldsymbol{\alpha}_{t}$ has bounded expectation for all $t$. This lemma is required both for proving the asymptotic convergence result in Theorem 1, which requires almost-sure boundedness of iterates and to obtain a better convergence rate in Theorem 2.
Lemma 6.
Under Assumptions 1-5,
$\mathbb{E}[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}]\leq C_{0},$
for some $C_{0}>0$, $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$ and all $t>0$.
The above three lemmas are sufficient for showing Theorem 1. We use the properties of two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$. This allows us to analyze the slower iteration, i.e., the control updates (8) in the limit $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$. We rewrite the iteration as a fixed point iteration with the mapping $g(\cdot)$. Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact Krasnosel’skiĭ–Mann algorithm [63]. We study its ODE limit and show convergence to the set of fixed points.
4.2 Finite Time Convergence Guarantees
In this subsection we detail how to obtain our finite time convergence guarantees. This result also uses the non-expansiveness on $g$, but requires a more nuanced argument. To prove the finite time mean square error bound, we bound the error as follows
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}]$ $\leq 2\mathbb{E}[\|\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))\|^{2}]$
$\;\;\;\;\;\;\;+2\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}].$
We first bound the mean square distance between $\boldsymbol{x}_{t}$ and $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, i.e., the NE corresponding to $\boldsymbol{\alpha}_{t}$.
Lemma 7.
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that
$\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t},$
for all $t>T_{0}$.
The key property used in the proof for this lemma is the strongly monotone property of the game which gives a recursive bound on $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$. This recursive bound gives a rate of $O(\eta_{t})$ when solved.
The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$. Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$. We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) Krasnosel’skiĭ–Mann algorithm. If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [64, 65]. Since this assumption is not possible in our case, we use a different approach and employ the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
Lemma 8.
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime\prime}>0$ and $T_{0}>0$ such that
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]\leq\frac{C^{\prime\prime}}{t\epsilon_{t}},$
for all $t>T_{0}$.
Lemma 7 and 8 together complete the proof for Theorem 2.
5 Simulation Results
(a) Loads for all resources
(b) Loads for 3 resources
Figure 2: Weighted Load Balancing in Resource Allocation Games
(a) Quadratic Cost
(b) Sum of rewards of all players
Figure 3: Quadratic Global Objective
In this section, we provide numerical simulations that include the two applications discussed in Section 2.1. For all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$. We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player. We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small. In each realization, $\left\{\alpha_{0}^{k}\right\}$ and $\left\{x_{n,0}^{k}\right\}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
For the resource allocation game (Figure 2), we design our game taking inspiration from DSM in electricity grid [66, 67]. We keep $N=1000$ and $d=24$. We choose the action set for each player $n$ as
$\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$
Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. The parameter $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$. We use the reward function
$r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$
Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{N}^{i}$, i.e., the weighted load at resource $i$. The coefficients $a_{n,i}^{(i)}$ are sampled independently from a Gaussian distribution with mean $1$ and variance $0.2$. The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are sampled independently from a Gaussian distribution with mean $35$ and variance $5$.
In Fig. 2(a), we plot the weighted load for each resource along with their target loads. We can observe that the system converges to a Nash equilibrium where the loads are balanced. In Fig. 2(b), we plot the weighted load for the first 3 resources. To study the effectiveness of our algorithm, we also plot the loads in the case of a uncontrolled system., i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$. It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
For the quadratic global objective (Figure 3), we consider a quadratic game. In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
We work with $N=100$ players and $d=5$, and action sets $\sum_{i=1}^{d}x_{n}^{i}\leq 10$ for each player $n$. To generate a random $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a Gaussian distribution with mean 0 and variance 1. We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$. A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$. The above matrices $\boldsymbol{Q_{n}}$ are obtained in a similar method as $\boldsymbol{G}$ but by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$. The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.
In Fig. 3(a), we observe that our algorithm is able to converge to a lower quadratic cost as compared to a system which is left uncontrolled. In Fig. 3(b), we plot the sum of rewards for our algorithm. This is compared with the sum of rewards if the global objective was directly optimized for, i.e., each player updates their action based on $\nabla\Phi(\boldsymbol{x})$ where $\Phi(\boldsymbol{x})$ is the quadratic cost. Our algorithm provides a much higher sum of rewards after convergence. The two plots together show that our algorithm can be used as a method to optimize for a global cost while still trying to maximize the rewards for players.
6 Conclusion
We studied how to control strongly monotone games where the game structure is unknown to the manager. The goal of the controller is to ensure the game satisfies target linear constraints at NE. Our algorithm only requires the manager to observe the instantaneous constraint violation. We prove that our simple online algorithm guarantees convergence to a point $\boldsymbol{x^{*}}$ which satisfies the linear constraints $\boldsymbol{A}\boldsymbol{x^{*}}=\boldsymbol{l}^{*}$ by adjusting control inputs $\boldsymbol{\alpha}$ in real-time. Further, we show that the convergence rate for our algorithm is $O(t^{-1/4})$. Our algorithm is a two time-scale algorithm, where the algorithm does not wait for the players to converge to a NE but uses the feedback from every turn. This is an encouraging result since our algorithm is easy to implement and maintains the users’ privacy. The manager does not observe individual actions, and players do not need to reveal their reward functions, which might even be unknown to them.
From a technical point of view, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$. Our algorithm essentially controls the unknown game such that its NE becomes desirable from an optimization point of view. Since convergence to NE is relatively fast, this has the potential to accelerate general cooperative multi-agent optimization [62].
Controlling unknown games with bandit feedback is a new online learning paradigm. The game to be controlled can be thought of as a non-stationary but highly structured bandit. Viewing game control as a bandit problem can result in both new exciting research questions, and new powerful game control techniques.
Generalizing our game control setting beyond linear constraints is an exciting new research avenue that can help tackle a wider class of problems.

### v2

@@
 Abstract
-Consider $N$ players each with a $d$-dimensional action set.
-Each of the players’ utility functions includes their reward function and a linear term for each dimension, with coefficients that are controlled by the manager.
-We assume that the game is strongly monotone, so if each player runs gradient descent, the dynamics converge to a unique Nash equilibrium (NE).
-The NE is typically inefficient in terms of global performance.
-The resulting global performance of the system can be improved by imposing $K$-dimensional linear constraints on the NE.
+Consider a game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager.
+We assume that the game is strongly monotone, so gradient play converges to a unique Nash equilibrium (NE).
+The NE is typically globally inefficient.
+The global performance at NE can be improved by imposing linear constraints on the NE.
 We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE.
-However, this requires knowing the players’ reward functions and their action sets.
-Obtaining this game structure information is infeasible in a large-scale network and violates the users’ privacy.
-To overcome this, we propose a simple algorithm that learns to shift the NE of the game to meet the linear constraints by adjusting the controlled coefficients online.
+However, this requires knowing the players’ reward functions and action sets.
+Obtaining this game information is infeasible in a large-scale network and violates user privacy.
+To overcome this, we propose a simple algorithm that learns to shift the NE to meet the linear constraints by adjusting the controlled coefficients online.
 Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets.
-We prove that our algorithm, which is based on two time-scale stochastic approximation, guarantees convergence with probability 1 to the set of NE that meet target linear constraints.
-We then provide a mean square convergence rate of $O(t^{-1/4})$ for our algorithm.
-This is the first such bound for two time-scale stochastic approximation where the slower time-scale is a fixed point iteration with a non-expansive mapping.
-We demonstrate how our scheme can be applied to optimizing a global quadratic cost at NE and load balancing in resource allocation games.
-We provide simulations of our algorithm for these scenarios.
+We prove that our algorithm converges with probability 1 to the set of NE that satisfy target linear constraints.
+We then prove an L2 convergence rate of near-$O(t^{-1/4})$.
 1 Introduction
@@
 However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy.
-As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5, 6, 7, 8, 9].
-Game theory is a natural way to tackle distributed decision-making [10, 11, 12].
+As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5].
+Game theory is a natural way to tackle distributed decision-making [6, 7].
 The game structure formalizes the effect of the interaction between users.
 The players may be cooperative, or they may be selfishly interested in maximizing their reward.
-In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they have no information regarding the reward functions of their peers.
+In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they do not know the reward functions of their peers.
 A Nash Equilibrium (NE) aims to predict the outcome of such a repeated interaction.
@@
 In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback.
-Such feedback is easy to monitor in practice and also maintains user privacy.
+Such feedback is easy to monitor in practice and maintains user privacy.
 Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
-We first prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1.
-We further show that our algorithm has a mean square error rate of $O(t^{-1/4})$ for the linear constraint.
-Due to the convex and compact action sets of the players, our slower time-scale is a fixed point iteration with a non-expansive mapping as opposed to the stronger contractive mapping assumed to analyze convergence rates in the stochastic approximation literature [16, 17].
-Our bound is the first of this kind for a two time-scale stochastic approximation scheme with a non-expansive mapping.
-Apart from tools from stochastic approximation, we employ variational inequalities and the Krasnosel’skiĭ–Mann algorithm in our analysis.
-Therefore, our simple algorithm offers a theoretically principled game control scheme for strongly monotone games.
+We prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1.
+We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraint.
+Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping typically assumed to analyze convergence rates [11, 12].
+Apart from stochastic approximation, we employ tools from variational inequalities and the Krasnosel’skiĭ–Mann (KM) algorithm in our analysis.
+Therefore, we provide a theoretically principled game control scheme for strongly monotone games.
 1.1 Related Work
-Distributed optimization [18, 19, 20] is concerned with optimizing a target objective distributedly over a network of agents.
+Distributed optimization [13, 14] is concerned with optimizing a target objective distributedly over a network of agents.
 However, these agents are non-interacting computational units so there is no game structure between them.
@@
 Moreover, distributed optimization entails significant communication overhead between agents.
-Distributed algorithms that converge to a NE have been studied in [21, 22, 23], including in monotone games [24, 25, 26].
-The players in our game use simple stochastic gradient play to converge to a NE.
-However, if the game is not controlled, the NE is typically inefficient from a global point of view.
+Distributed algorithms that converge to NE have been studied in [15, 16, 17], including in monotone games [18, 19, 20].
+The players in our game use simple stochastic gradient play to converge to NE.
+However, if the game is uncontrolled, the NE is typically globally inefficient.
 Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
-Compared to mechanism design approaches [27, 28, 29], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only receives the sum of actions (loads) of all players.
-Our approach resembles the idea of a Stackelberg game [30, 31, 32, 33], where the manager is the leader and the players are the followers (usually a single follower is considered).
-However, game-theoretic algorithms converge to a NE that can be inefficient from an optimization point of view.
+Compared to mechanism design approaches [21, 22], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only observes the constraint violation.
+Our approach resembles a Stackelberg game [23, 24, 25], where the manager is the leader and the players are the followers (usually a single follower is considered).
+However, game-theoretic algorithms converge to a NE that can be globally inefficient.
 Our work provides a mechanism that provably shifts the unique NE to a point that satisfies target linear constraints that can force the NE to be efficient (e.g., first-order optimality condition for a global quadratic objective).
@@
 Instead, the rewards of this bandit are generated by the game dynamics.
-This “game bandit” also introduces a new type of noise.
+This “game bandit” introduces a new type of noise.
 The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints.
-Hence, the type of feedback that the manager needs is about the behavior at NE.
+Hence, the type of feedback the manager needs concerns the behavior at NE.
 However, the dynamics are not at NE every turn.
@@
 Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
-The most closely related literature to our work is that on intervention and control of games [36, 37, 38, 39, 40, 41, 42, 43, 44].
+Our problem can be viewed as a Generalized Nash Equilibrium Problem (GNEP), where players seek to converge to a NE subject to coupled constraints [28].
+Efficient computation of GNE in strongly monotone games in the noiseless setting has been widely studied [29, 30, 31].
+Stochastic convergence to GNEs has recently been studied under different assumptions.
+In [32] players share their actions with others, and in [33] the reward function is assumed to be linear.
+In [34], stochastic convergence to GNE is studied more generally, resulting in a centralized algorithm where the weighted empirical average of the actions is shown to converge in expectation.
+We provide a distributed stochastic learning scheme that converges in probability 1 to GNE with linear coupled constraints in strongly monotone games without communication between players.
+Our work belongs to the literature on intervention and control of games [35, 36, 37, 38, 39, 40, 41, 42, 43].
 In [44] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE.
@@
 The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets.
-For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., their geometric locations).
+For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., geometric locations).
 Additionally, the pricing scheme in [46, 47, 48] requires the manager to know the reward functions.
 Control of dynamic Markov potential games was considered in [40], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
-There have been works obtaining convergence rate for contractive two time-scale stochastic approximation [50, 16].
+Some works obtained convergence rates for contractive two time-scale stochastic approximation [49, 11].
 These works assume that both the time-scales are fixed point iterations with contractive mappings.
 We consider a more general case where the slower timescale has a non-expansive mapping.
-A preliminary conference version, limited to load balancing in resource allocation, was published in [51].
-In this paper, we extend the framework in [51] to linear constraint in strongly monotone games.
-In particular, the NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$.
+An earlier conference paper [50] was limited to load balancing in resource allocation.
+This paper’s framework is similar to that in [50], but is extended to linear constraint in strongly monotone games.
+The NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$.
+Nevertheless, the results here are completely novel and independent of [50].
 Our convergence guarantees are with probability 1 instead of in L2 like in [50].
-Last but not least, we provide a finite time convergence error bound.
-Our novel results require new analysis based on non-expansive mappings that has limited overlap with the analysis in [51].
-1.2 Outline
+Unlike [50], we provide a finite time convergence error bound.
+Our novel analysis is based on two time-scale stochastic approximation with non-expansive mappings.
+1.2 Outline and Notation
 The rest of the paper is organized as follows.
-In Section 2 we formulate our game control problem and detail our assumptions.
-In Subsection 2.1 we detail two applications that fall under our general model.
-In Section 3 we describe the algorithm that the manager uses to control the game and state our main results regarding its performance guarantees.
-In Section 4 we break the proofs of our main results into lemmas.
-In Section 5 we provide numerical simulations for the application described earlier.
-Then, In Section 6 we conclude the paper and propose future research directions.
-1.3 Notation
+Section 2 formulates our game control problem and details our assumptions.
+Subsection 2.1 details two applications that fall under our general model.
+Section 3 describes the game control algorithm and states our main results regarding its performance guarantees.
+Section 4 breaks the proofs of our main results into lemmas.
+Section 5 provides numerical simulations for the application described earlier.
+Finally, Section 6 concludes the paper and proposes future research directions.
 […]
@@
 We use $\boldsymbol{x}\in\mathcal{X}$ to denote the concatenation of all action vectors.
-Then the utility for each player is defined as:
-$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{d}x_{n}^{d}.$ (1)
+The utility for each player is defined as:
+$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
 Here $r_{n}(\boldsymbol{x})$ is the reward function for player $n$.
@@
 $\|F(\boldsymbol{y})-F(\boldsymbol{x})\|\leq L\|\boldsymbol{y}-\boldsymbol{x}\|.$ (4)
-Our game evolves in discrete time, with players taking actions $\boldsymbol{x}_{n,t}$, given the controlled coefficients $\boldsymbol{\beta}_{n,t}$, at time $t$.
-The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the objective of the game manager (discussed below (6)).
-Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., the game is uncontrolled), then the players will eventually converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
+Our game evolves in discrete time, so at time $t$ players take actions $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$.
+The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the manager’s objective (discussed below (6)).
+Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
 In many scenarios, the players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive (i.e., bandit feedback) or other data they collect.
@@
 Then $\boldsymbol{g}_{t-1}=F(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{t}$, where $\boldsymbol{M}_{t}$ satisfies
-$\mathbb{E}[\boldsymbol{M}_{t}\mid\mathcal{F}_{t-1}]=0\;\;\;\textrm{and}\;\;\;\mathbb{E}[\|\boldsymbol{M}_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq M_{c}.$ (5)
-Here $\mathcal{F}_{t}$ is the filtration that summarizes the past, i.e.,
+$\mathbb{E}[\boldsymbol{M}_{t}\mid\mathcal{F}_{t-1}]=0\;\;\;\textrm{and}\;\;\;\mathbb{E}[\|\boldsymbol{M}_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq M_{c},$ (5)
+for some $M_{c}>0$ and a filtration $\mathcal{F}_{t}$ that summarizes the past:
 $\mathcal{F}_{t}=\sigma(\boldsymbol{x}_{\tau},\boldsymbol{\beta}_{\tau}\mid\forall\tau\leq t).$
@@
 2.1.1 Quadratic Global Objective
-For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global objective of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$ which they wish to minimize.
+For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global cost of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$.
 All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$.
-Hence, using our approach, we can set $A=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a local optimum.
+Hence, using our approach, we can set $\boldsymbol{A}=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a global optimum.
 2.1.2 Weighted Resource Allocation Games
@@
 Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol.
-In the electricity grid, the manager can be the utility company running DSM to cut the production costs of energy at peak hours.
+In the electricity grid, the manager can be the utility company running Demand Side Management (DSM) to cut the production costs of energy at peak hours.
 In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network.
@@
 In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources.
-We have $d$ constraints i.e., $K=d$ with the idea of each linear constraints being to meet the target loads for each resource $i$, i.e.,
+We have $d$ linear constraints i.e., $K=d$ and each represents a target load for each resource $i$, i.e.,
 […]
-The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize for application-specific objectives such as system efficiency, regulations, and operation costs.
-The controlled coefficients $\boldsymbol{\beta}$ can be thought of as incentives to use a resource in this application.
+The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize application-specific objectives such as system efficiency, regulations, and operation costs.
+The controlled coefficients $\boldsymbol{\beta}$ represent incentives to use a resource.
 The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
-A special case of this is resource allocation without weights, studied in [51]. In this case the linear constraints are
-[…]
-Generalizing to a weighted sum of actions allows modeling of cases such as power control in wireless networks and load balancing in data-centers.
+A special case is resource allocation without weights, studied in [50].
+Generalizing to a weighted sum of actions allows modeling of power control in wireless networks and load balancing in data centers.
 In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains.
-In data-centers, different users send jobs of different sizes to the different servers (resources), so their actions (how many jobs to send where) affect the loads differently.
+In data centers, different users send jobs of different sizes to different servers (resources), so the load is a weighted sum of their actions (how many jobs to send where).
 3 Game Control Algorithm
@@
 Initialization: Let $\boldsymbol{x}_{0}\in\mathcal{X}$ and $\boldsymbol{\alpha}_{0}\in\mathbb{R}^{K}$.
-Input: Target vector $\boldsymbol{l}^{*}$.
+Input: Target vector $\boldsymbol{\ell^{*}}$.
 For each turn $t\geq 1$ do
@@
 The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$.
-To overcome this, we propose an online learning approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$.
+To overcome this, we propose an online approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$.
 We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$.
 This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
-Our algorithm can be thought of as a two time-scale stochastic approximation algorithm.
+Our algorithm is a two time-scale stochastic approximation algorithm.
 In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$.
 In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
-This algorithm can also be thought of as an iterative approach to solving the constrained optimization problem using Lagrange multipliers.
+Essentially, our algorithm is an iterative approach to solving the constrained optimization problem using Lagrange multipliers.
 We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
-The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied at all times as that will require communication between agents, and hence the manager uses iterative approach to solve the constrained optimization problem.
+The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied for all $t$ without communicating, and hence the manager uses an iterative approach instead.
 Similar approaches have been used in [57, 58] for different problems.
@@
 This behavior can be encountered with selfish players in a large-scale network or with cooperative players.
-In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [60, 61, 15].
+In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [59, 10].
 In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps.
@@
 The manager’s algorithm in (8) is robust against strategic manipulations of a small group of players since each player has a negligible effect on $\boldsymbol{A}\boldsymbol{x}_{t-1}$, which is its only way to impact (8).
-It is not clear if even a large group of players can manipulate (8) to increase their accumulated reward over time since any manipulated change to $\boldsymbol{\alpha}_{t}$ would be temporary and incurs losses to the players.
-Moreover, truthful reporting is not an issue in our algorithm since the players do not report anything, and the manager cannot observe their actions.
-Instead, the manager only needs to measure the instantaneous constraint violation
+We conjecture that even a large group of players cannot manipulate (8) to increase their accumulated reward over time since any manipulated change to $\boldsymbol{\alpha}_{t}$ would be temporary and incurs losses to the players.
+Moreover, truthful reporting is not an issue since the players do not report anything, and the manager cannot observe their actions.
+Instead, the manager only needs to measure the instantaneous constraint violation.
 3.1 Performance Guarantees
@@
 where $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ denotes the NE corresponding to $\boldsymbol{\alpha}$.
-We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is is just Slater’s condition [63] for convex optimization.
+We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is just Slater’s condition [61] for convex optimization.
 Assumption 4.
-The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is a continuously differentiable convex function and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
-We also take the following standard assumptions on our stepsize sequences [16].
+The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
+We also make the following standard assumptions on our step-size sequences [11].
 Assumption 5.
-The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ which satisfy:
+The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ which satisfy:
   1. (a)
@@
 Theorem 1.
-Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and
+Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and
 $\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}},$
@@
 $\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right).$
-Therefore the best rate is obtained for $\eta=0.5$ and $\epsilon=0.75$:
-$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-1/4}\right).$
+Therefore the best rate is obtained for $\eta=0.5+\delta/3$ and $\epsilon=0.75+\delta$:
+$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right),$
+where $\delta$ is arbitrarily small.
 The above theorem holds for all $t$ if we choose $T_{1}$ and $T_{2}$ to be sufficiently large and then $\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right)$ for all $t$, for some $C>0$.
+A smaller value of $\epsilon$ gives a better rate.
+The optimal rate is then dictated by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\uparrow\infty}\epsilon_{t}^{2}/\eta_{t}^{3}=0$ ($2\epsilon>3\eta$).
 4 Convergence Analysis
-In this section, we detail our proof strategy for Theorem 1 and 2. The proofs for the lemmas below and the main theorems are given in Appendix I and II respectively.
-Throughout the proofs, we use the notation $F(\boldsymbol{x},\boldsymbol{\alpha})=F(\boldsymbol{x})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$, which is the gradients vector of the game given the controlled input $\boldsymbol{\alpha}$.
-Our first lemma shows that the desired set $\mathcal{N}_{opt}$ is non-empty. The proof for this lemma follows from the strongly monotone nature of the game and Slater’s conditions (Assumption 4).
+In this section, we detail our proof strategy for Theorem 1 and 2.
+Throughout the proofs, we denote the gradient vector given the controlled input $\boldsymbol{\alpha}$ by $F(\boldsymbol{x},\boldsymbol{\alpha})=F(\boldsymbol{x})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$.
+Our first lemma shows that the desired set $\mathcal{N}_{opt}$ is non-empty. The proof follows from the strong monotonicity of the game and Slater’s conditions (Assumption 4).
 Lemma 3.
 Under Assumptions 1 and 4, there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, i.e., the set $\mathcal{N}_{opt}$ is non-empty.
+Proof.
+Let $\mathcal{G}(\boldsymbol{\alpha})$ be our $N$-player game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and action set $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$.
+Consider a modified $N$-player game $\mathcal{G}^{\prime}$ with utility function $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ for $n\in\{1,\ldots,N\}$ and action set $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$.
+Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 along with equations (3.3) and (3.6) from [8] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$, and $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
+$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
+and
+$\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}\;\textrm{and}\;h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\}.$
+Let $\boldsymbol{\alpha}=-\boldsymbol{\lambda}$. Then the above equations can be written as
+$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}=0$
+and
+$h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\},$
+which implies that $\hat{\boldsymbol{x}}$ is a NE for the game $\mathcal{G}(\boldsymbol{\alpha})$, and hence $\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\hat{\boldsymbol{x}}$.
+We also know that $\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}$.
+Hence there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, showing that $\mathcal{N}_{opt}$ is non-empty.
 One of the main technical challenges in our analysis stems from the constrained actions represented by the convex and compact action sets.
@@
     where $\mu$ is the monotonicity parameter of the game.
+Proof.
+For a fixed $\boldsymbol{\alpha}$, $-F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively.
+Then [51, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
+$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$
+Hence for $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$, we have
+$\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|$
+$\leq L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
+=$\stackrel{{\scriptstyle(a)}}{{=}}L^{\prime}\Big{\|}\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
+$\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\Big{\|}$
+$\leq L^{\prime}\|F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})-F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1})\|$
+$=L^{\prime}\|\boldsymbol{A}^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})\|$
+$\leq L_{0}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|,$
+where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$. Here equality (a) follows from [51, Proposition 1.5.8] which says
+$\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})=\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})).$
+This completes the proof for part (a) of Lemma 4.
+Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$ Since $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})$ is the NE corresponding to $\boldsymbol{\alpha}_{1}$, using [51, Proposition 1.4.2], we have for every $\boldsymbol{x}\in\mathcal{X}$ that $\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0$.
+For $\boldsymbol{x}=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$
+$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0.$
+Similarly, we also have
+$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)\right\rangle\leq 0.$
+By combining the above two inequalities we obtain
+$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\geq 0.$ (11)
+Then
+$-\mu\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
+$\stackrel{{\scriptstyle(a)}}{{\geq}}\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\rangle$
+$=\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\rangle$
+$\;\;\;+\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{2}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{1}\rangle$
+$\stackrel{{\scriptstyle(b)}}{{\geq}}\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{2}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{1}\rangle$
+$=(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}))^{\intercal}\boldsymbol{A}^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})$
+$=(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}))^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})$
+$=\left\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\right\rangle.$
+Here inequality (a) uses that $F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone in $\boldsymbol{x}$ with parameter $\mu$ and (b) uses (11). ∎
 The next lemma shows that the map $g(\cdot)$ is non-expansive.
@@
 i.e., the map $g(\cdot)$ is non-expansive.
+Proof.
+For any $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$,
+$\|g(\boldsymbol{\alpha}_{2})-g(\boldsymbol{\alpha}_{1})\|^{2}$
+$=\|\boldsymbol{\alpha}_{2}+\gamma\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{\alpha}_{1}-\gamma\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
+$=\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}+\gamma^{2}\left\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\right\|^{2}$
+$\;\;\;+2\gamma\left\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\right\rangle$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}+\gamma^{2}\|\boldsymbol{A}\|^{2}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
+$\;\;\;-2\mu\gamma\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
+$\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}.$
+Here inequality (a) follows from part (b) of Lemma 4 and inequality (b) follows from the assumption on $\gamma$ that $2\mu\gamma>\gamma^{2}\|\boldsymbol{A}\|^{2}$. Hence $g(\cdot)$ is non-expansive. ∎
 The next lemma shows that the iterate $\boldsymbol{\alpha}_{t}$ has bounded expectation for all $t$.
@@
 for some $C_{0}>0$, $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$ and all $t>0$.
-The above three lemmas are sufficient for showing Theorem 1.
+Proof.
+Let $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$.
+We first give a recursive bound on the term $\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$.
+For this, note that
+$\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
+$=\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
+=$\stackrel{{\scriptstyle(a)}}{{=}}\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1}))\|^{2}$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}\|\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
+$=\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1})$
+$\;\;+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
+$=\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$ (13a)
+$\;\;+\|\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$ (13b)
+$\;\;+2\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}.$ (13c)
+Here equality (a) uses that $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\in\mathcal{X}$ and inequality (b) follows from since the projection operator $\Pi_{\mathcal{X}}(\cdot)$ for any convex set $\mathcal{X}$ is non-expansive.
+We first simplify the term (13a):
+$\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$
+$=\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\eta_{t}^{2}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$
+$\;\;+2\eta_{t}\langle\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\rangle.$
+The last term is bounded as follows.
+$\langle\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\rangle$
+$\leq\langle\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})-F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})\rangle$
+$\leq-\mu\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}.$
+Here the first inequality follows since $\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}),F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}),\boldsymbol{\alpha})\rangle\leq 0$, for all $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ and $\boldsymbol{x}\in\mathcal{X}$, since $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is a NE (Proposition 1.4.2 in [51]) and the second inequality follows from the strong monotonicity property of the game.
+We also have that
+$\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}=\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}\|^{2}$
+$=\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}\|^{2}$
+$\leq 2\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}\|^{2}+2\|\boldsymbol{A}^{\intercal}\|^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}$
+$\leq c_{1}^{2}+c_{2}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}.$
+where the last inequality follows for some $c_{1}$ and $c_{2}$ as $\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}\|$ is bounded by some constant as $\boldsymbol{x}_{t}\in\mathcal{X}$, where $\mathcal{X}$ is compact.
+Coming back to (13a), we have
+$\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$
+$\leq(1-2\mu\eta_{t})\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\eta_{t}^{2}c_{1}^{2}+\eta_{t}^{2}c_{2}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}.$
+For (13b), note that
+$\mathbb{E}[\|\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}\mid\mathcal{F}_{t}]$
+$=\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\mathbb{E}[\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}\mid\mathcal{F}_{t}]$
+$\;\;+2\eta_{t}\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]$
+=$\stackrel{{\scriptstyle(a)}}{{=}}\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}\eta_{t}^{2}M_{c}+L_{0}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha}_{t+1}\|^{2}=\eta_{t}^{2}M_{c}+L_{0}^{2}\epsilon_{t}^{2}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|^{2}$
+$\stackrel{{\scriptstyle(c)}}{{\leq}}\eta_{t}^{2}M_{c}+\epsilon_{t}^{2}c_{3}^{2}.$
+Here equality (a) follows from the fact that $\alpha_{t+1}$ is $\mathcal{F}_{t}$ measurable and $\boldsymbol{M}_{t+1}$ is a Martingale difference sequence, hence making the term $\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]=0$.
+Inequality (b) follows from Lemma 4 part (a), and (c) follows from $\boldsymbol{x}_{t}\in\mathcal{X}$ for some constant $c_{3}$.
+Finally, for the third term (13c):
+$2\mathbb{E}\big{[}\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}\mid\mathcal{F}_{t}\big{]}$
+$=2\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|L_{0}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha}_{t+1}\|$
+$\stackrel{{\scriptstyle}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|$
+$\;\;+2\eta_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}c_{3}+2\eta_{t}\epsilon_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|c_{3}.$
+Here the first equality uses that $\boldsymbol{M}_{t+1}$ is a martingale difference sequence and $\boldsymbol{x}_{t}$, $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{t+1}$ are $\mathcal{F}_{t}$ measurable.
+Inequality (a) follows from the Cauchy-Schwarz inequality and the Lipschitz continuity of the map $\boldsymbol{x^{*}}(\cdot)$ (Lemma 4) and (b) uses boundedness of $\boldsymbol{x}_{t}$ and similar to previous expansions.
+To bound the term $2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}c_{3}$, we use $(2ab\leq a^{2}+b^{2})$ where $a=\sqrt{\mu\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}/2}$ and $b=\sqrt{(\epsilon_{t}^{2}/\eta_{t})(2c_{3}^{2}/\mu)}$.
+To bound the term $2\eta_{t}\epsilon_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|c_{3}$, we use the same inequality with $a=\eta_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|$ and $b=\epsilon_{t}c_{3}$.
+This yields
+$2\mathbb{E}\big{[}\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}\mid\mathcal{F}_{t}\big{]}$
+$\leq\frac{\mu}{2}\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu+\eta_{t}^{2}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}+\epsilon_{t}^{2}c_{3}^{2}$
+$\leq\frac{\mu}{2}\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu$
+$\;\;\;+\eta_{t}^{2}c_{1}^{2}+\eta_{t}^{2}c_{2}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\epsilon_{t}^{2}c_{3}^{2}$
+We have now bounded all three terms from (13). Combining them after taking expectation gives us:
+$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$
+$\leq(1-\frac{3}{2}\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+\eta_{t}^{2}(2c_{1}^{2}+M_{c})$
+$\;\;\;+2\epsilon_{t}^{2}c_{3}^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right].$
+Our assumption that $\lim_{t\to\infty}\epsilon_{t}^{2}/\eta_{t}^{3}\rightarrow 0$ implies that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for some $C_{1}>0$ and $t>T_{0}$. Then for all $t>T_{0}$:
+$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$ $\leq(1-1.5\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$
+$+$ $c_{4}\eta_{t}^{2}+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right],$ (14)
+for $c_{4}=2c_{1}^{2}+M_{c}+2C_{1}c_{3}^{2}+2C_{1}c_{3}^{2}/\mu$.
+We next give a recursive bound on $\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$. Let $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$. Then iteration (8) can be written as
+$\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t})+\epsilon_{t}^{\prime}\boldsymbol{A}\gamma(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})),$
+where $g(\boldsymbol{\alpha})=\boldsymbol{\alpha}+\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$.
+For ease of notation define the sequences $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}$ and $q_{t}=\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))$.
+Let $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$ be a fixed point of $g(\cdot)$.
+Then
+$\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}$
+$=\|\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}+q_{t})-\boldsymbol{\alpha^{*}}\|^{2}$
+$=\|(1-\epsilon_{t}^{\prime})(\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}})+\epsilon^{\prime}_{t}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}+q_{t})\|^{2}.$
+Using Corollary 2.14 from [61], we obtain
+$\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}$ $\leq(1-\epsilon_{t}^{\prime})\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}$
+$\;\;\;+\epsilon_{t}^{\prime}\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}+q_{t}\|^{2}$
+$\;\;\;-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\|p_{t}+q_{t}\|^{2}.$ (15)
+We bound the second term as follows:
+$\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}+q_{t}\|^{2}$
+$=\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}\|^{2}+\|q_{t}\|^{2}+2\langle g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}},q_{t}\rangle$
+$\leq\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\|q_{t}\|^{2}+2\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}\|\|q_{t}\|$
+$\leq\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\|q_{t}\|^{2}+2\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|\|q_{t}\|$
+Here the first inequality stems from the non-expansive nature of map $g(\cdot)$ and the final inequality uses the fact that $g(\boldsymbol{\alpha^{*}})=\boldsymbol{\alpha^{*}}$ along with the non-expansive nature of $g(\cdot)$.
+Now
+$2\epsilon_{t}^{\prime}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|\|q_{t}\|\leq 2\epsilon_{t}\|\boldsymbol{A}\|\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|$
+$\leq\mu\eta_{t}/2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{2\|\boldsymbol{A}\|^{2}\epsilon_{t}^{2}}{\mu\eta_{t}}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}.$
+Then returning to (4.1),
+$\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\leq\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\epsilon_{t}^{\prime}\gamma^{2}\|\boldsymbol{A}\|^{2}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}$
+$\;\;\;+\mu\eta_{t}/2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{2\|\boldsymbol{A}\|^{2}\epsilon_{t}^{2}}{\mu\eta_{t}}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}$
+$\;\;\;-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\|p_{t}+q_{t}\|^{2}.$
+Again using our assumption that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for some appropriate $C_{1}>0$ and $t>T_{0}$, and taking expectation we have
+$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
+$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\mu\eta_{t}\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}\right]$
+$\;\;\;+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\mathbb{E}[\|p_{t}+q_{t}\|^{2}].$ (16)
+Adding (4.1) and (4.1), we get
+$\mathbb{E}\left[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
+$\leq(1+4c_{2}^{2}\eta_{t}^{2})\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+c_{4}\eta_{t}^{2}.$
+Expanding the recursion for $t>T_{0}$ gives us
+$\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
+$\leq\mathbb{E}\left[\|\boldsymbol{x}_{T_{0}}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{T_{0}})\|^{2}+\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}\right]\prod_{i=T_{0}}^{t-1}(1+4c_{2}^{2}\eta_{i}^{2})$
+$\;\;\;+c_{4}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}\prod_{j=i+1}^{t-1}(1+4c_{2}^{2}\eta_{j}^{2})$
+$\leq\mathbb{E}\left[\|\boldsymbol{x}_{T_{0}}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{T_{0}})\|^{2}+\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}\right]e^{\left(4c_{2}^{2}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}\right)}$
+$\;\;\;+c_{4}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}e^{\left(4c_{2}^{2}\sum_{j=i+1}^{t-1}\eta_{j}^{2}\right)}.$
+Here the final inequality uses that $(1+x)\leq e^{x}$ for any $x\in\mathbb{R}$.
+Now we use our assumption that $\eta_{t}$ is square-summable.
+Hence, $\exp(4c_{2}^{2}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2})$ and $\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}\exp\left(4c_{2}^{2}\sum_{j=i+1}^{t-1}\eta_{j}^{2}\right)$ are bounded by constants.
+Finally, $T_{0}$ is a constant dependent only on the stepsize sequence, and hence $\|\boldsymbol{x}_{T_{0}}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{T_{0}})\|^{2}+\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}$ can be bounded by a constant using the discrete Gronwall’s inequality [62].
+This finally gives us that for some constant $C_{0}$
+$\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]\leq C_{0},$
+which completes the proof for Lemma 6. ∎
+The above three lemmas are sufficient for proving Theorem 1 (see Appendix).
 We use the properties of two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$.
@@
 We rewrite the iteration as a fixed point iteration with the mapping $g(\cdot)$.
-Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact Krasnosel’skiĭ–Mann algorithm [63].
+Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact KM algorithm [61].
 We study its ODE limit and show convergence to the set of fixed points.
@@
 In this subsection we detail how to obtain our finite time convergence guarantees.
-This result also uses the non-expansiveness on $g$, but requires a more nuanced argument.
-To prove the finite time mean square error bound, we bound the error as follows
+This result also uses the non-expansiveness on $g$, but requires a more nuanced argument that bounds the error as follows
 $\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}]$ $\leq 2\mathbb{E}[\|\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))\|^{2}]$
@@
 We first bound the mean square distance between $\boldsymbol{x}_{t}$ and $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, i.e., the NE corresponding to $\boldsymbol{\alpha}_{t}$.
+The key to proving this lemma is the strong monotonicity of the game which gives a recursive bound on $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$.
+This recursive bound gives a rate of $O(\eta_{t})$ when solved.
 Lemma 7.
-Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that
+Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
+$\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t}.$
+Proof.
+Returning to the the recursive bound (4.1) from the proof for Lemma 6, and using Lemma 6 that shows that $\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]\leq C_{0}$, we now obtain the recursion
+$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$ $\leq(1-1.5\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$
+$+\eta_{t}^{2}(c_{4}+2c_{2}^{2}C_{0}).$
+For $\eta_{t}=(t+T_{1})^{-\eta}$ with $10<\eta<$ 1, this recursion yields [12]
 $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t},$
-for all $t>T_{0}$.
-The key property used in the proof for this lemma is the strongly monotone property of the game which gives a recursive bound on $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$.
-This recursive bound gives a rate of $O(\eta_{t})$ when solved.
+for $t>T_{0}$ for some constant $C^{\prime}$. ∎
 The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$.
 Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$.
-We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) Krasnosel’skiĭ–Mann algorithm.
-If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [64, 65].
-Since this assumption is not possible in our case, we use a different approach and employ the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
+We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) KM algorithm.
+If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [63].
+Since this assumption is not true in our case, we use the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
 Lemma 8.
-Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime\prime}>0$ and $T_{0}>0$ such that
-$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]\leq\frac{C^{\prime\prime}}{t\epsilon_{t}},$
-for all $t>T_{0}$.
-Lemma 7 and 8 together complete the proof for Theorem 2.
+Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime\prime}>0$ and $T_{0}>0$ such that for all $t>T_{0}$.
+$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]\leq\frac{C^{\prime\prime}}{t\epsilon_{t}}.$
+Proof.
+Define $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}$, $q_{t}=\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))$ and $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$.
+Returning to the recursive bound (4.1) from the proof of Lemma 6, and now using Lemma 6 and Lemma 7,
+$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
+$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\eta_{t}^{2}c_{5}^{2}-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\mathbb{E}[\|p_{t}+q_{t}\|^{2}]$ (17)
+for some constant $c_{5}$. For the third term above, we note that
+$-\|p_{t}+q_{t}\|^{2}$ $=-\|p_{t}\|^{2}-\|q_{t}\|^{2}-2\langle p_{t},q_{t}\rangle$
+$\leq-\|p_{t}\|^{2}-\|q_{t}\|^{2}+2\|p_{t}\|\|q_{t}\|$
+$\leq-\|p_{t}\|^{2}+c_{6}\|q_{t}\|.$
+Note that $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}})$.
+Hence the last inequality follows from the compactness of $\mathcal{X}$ for some constant $c_{6}>0$.
+This implies $\mathbb{E}\left[-\|p_{t}+q_{t}\|^{2}\right]\leq\mathbb{E}\left[-\|p_{t}\|^{2}\right]+2c_{7}\sqrt{\eta_{t}},$ for some constant $c_{7}>0$.
+This inequality follows from Lemma 7 and since $\mathbb{E}[\|q_{t}\|]^{2}\leq\mathbb{E}[\|q_{t}\|^{2}]$.
+Using our assumption that $\epsilon_{t}$ goes to zero, we have that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$.
+For the rest of the proof, we pick $T_{0}$ to be large enough such that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$, $\epsilon_{t}<\gamma/2$ and $\eta_{t}<C_{2}$ for all $t>T_{0}$, for appropriate $C_{1}$ and $C_{2}$.
+Returning to (4.2), for all $t>T_{0}$, we have
+$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
+$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\eta_{t}^{2}c_{5}^{2}+\epsilon_{t}^{\prime}\sqrt{\eta_{t}}c_{7}-(\epsilon_{t}^{\prime}/2)\mathbb{E}[\|p_{t}\|^{2}].$
+This gives us
+$(\epsilon_{t}^{\prime}/2)\mathbb{E}[\|p_{t}\|^{2}]$ $\leq\mathbb{E}[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}]-\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$
+$\;\;\;+\eta_{t}^{2}c_{5}^{2}+\epsilon_{t}^{\prime}\sqrt{\eta_{t}}c_{7}.$
+Summing from $i=T_{0}$ to $t$ gives us
+$\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\mathbb{E}[\|p_{i}\|^{2}]$ $\leq\mathbb{E}[\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}]-\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$
+$\;\;+\sum_{i=T_{0}}^{t}(\eta_{i}^{2}c_{5}^{2}+\epsilon_{i}^{\prime}\sqrt{\eta_{i}}c_{7}).$ (18)
+The left hand term in the above inequality depends on $\|p_{i}\|^{2}$ for all $i\in\{T_{0},\ldots,t\}$.
+We next show a bound for $\|p_{i}\|^{2}$ in terms of $\|p_{t}\|^{2}$ for all $i\in\{T_{0},\ldots,t\}$.
+For this, we first recall that $\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(p_{t}+q_{t}).$ Next,
+$\|p_{t+1}\|^{2}=\|p_{t}\|^{2}+\|p_{t+1}-p_{t}\|^{2}+2\langle p_{t+1}-p_{t},p_{t}\rangle$
+$\leq\|p_{t}\|^{2}+\|p_{t+1}-p_{t}\|^{2}+\frac{2}{\epsilon_{t}^{\prime}}\langle p_{t+1}-p_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}-\epsilon_{t}^{\prime}q_{t}\rangle$
+Now,
+$2\langle p_{t+1}-p_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle$
+$=2\langle g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t+1}+\boldsymbol{\alpha}_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle$
+$=2\langle g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle-2\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}$
+$=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-2\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}$
+$\;\;-\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t+1}+\boldsymbol{\alpha}_{t}\|^{2}$
+$=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}-\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-\|p_{t+1}-p_{t}\|^{2}$
+$\leq-\|p_{t+1}-p_{t}\|^{2}.$
+Here the final inequality follows from the non-expansive nature of $g(\cdot)$. This gives us
+$\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\|p_{t+1}-p_{t}\|^{2}-2\langle p_{t+1}-p_{t},q_{t}\rangle$
+$=\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\left\|p_{t+1}-p_{t}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}q_{t}\right\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}$
+$\leq\|p_{t}\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}.$
+Using the assumption that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
+$\mathbb{E}[\|p_{i}\|^{2}]$ $\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2\epsilon_{j}^{\prime}\mathbb{E}\|q_{j}\|^{2}$
+$\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2c_{8}\epsilon_{j}^{\prime}\eta_{j},$
+for some constant $c_{8}>0$. Returning to (4.2), we now have
+$\mathbb{E}[\|p_{t}\|^{2}]\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\leq\mathbb{E}[\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}]$
+$\;\;+\sum_{i=T_{0}}^{t}(\eta_{i}^{2}c_{5}^{2}+\epsilon_{i}^{\prime}\sqrt{\eta_{i}}c_{7})+2c_{8}\sum_{i=T_{0}}^{t}\epsilon_{i}^{\prime}\sum_{j=i}^{t-1}\epsilon_{j}^{\prime}\eta_{j}.$ (19)
+Now we use our assumption that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for all $t>T_{0}$.
+Then $\epsilon_{i}^{\prime}\sqrt{\eta_{i}}=O(\eta_{i}^{2})$ and hence the sequence $\{\epsilon_{i}^{\prime}\sqrt{\eta_{i}}\}$ is summable.
+Finally for the double summation, we use the integral test and Assumption 5 to obtain
+$\sum_{i=T_{0}}^{t}\epsilon_{i}^{\prime}\sum_{j=i}^{t-1}\epsilon_{j}^{\prime}\eta_{j}\leq\frac{1}{\gamma^{2}}\sum_{i=T_{0}}^{\infty}\frac{1}{(i+T_{2})^{\epsilon}}\sum_{j=i}^{\infty}\frac{1}{(j+T_{3})^{\epsilon+\eta}}$
+$\leq\sum_{i=T_{0}}^{\infty}\frac{1}{(i+T_{2})^{\epsilon}}\frac{c_{9}}{(i+T_{3})^{\epsilon+\eta-1}}\leq\sum_{i=T_{0}}^{\infty}\frac{c_{10}}{(i+T_{3})^{4\eta-1}}.$
+Here $c_{9}$ and $c_{10}$ are constants and $T_{3}=\max\{T_{1},T_{2}\}$.
+The final inequality follows from the fact that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ and hence $2\epsilon\geq 3\eta$.
+Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $14\eta-1>$ 1, so the summation is bounded by a constant.
+Hence all terms in the right-hand side of (4.2) are bounded by a constant.
+Then for some $c_{11}>0$, we have
+$\mathbb{E}[\|p_{t}\|^{2}]\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\leq c_{11}.$
+Now $\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\geq(t-T_{0}+1)\epsilon_{t}^{\prime}/2$.
+This finally gives us the bound $\mathbb{E}[\|p_{t}\|^{2}]\leq C^{\prime\prime}/(t\epsilon_{t}),$ for all $t>T_{0}$.
+∎
 5 Simulation Results
@@
 Figure 3: Quadratic Global Objective
-In this section, we provide numerical simulations that include the two applications discussed in Section 2.1.
-For all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$.
+In this section, we simulate the two applications discussed in Section 2.1.
+In all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$.
 We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player.
 We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small.
-In each realization, $\left\{\alpha_{0}^{k}\right\}$ and $\left\{x_{n,0}^{k}\right\}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
-For the resource allocation game (Figure 2), we design our game taking inspiration from DSM in electricity grid [66, 67]. We keep $N=1000$ and $d=24$. We choose the action set for each player $n$ as
-$\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$
-Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$.
-The parameter $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$.
-We use the reward function
-$r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$
-Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$.
+In each realization, $\alpha_{0}^{k}$ and $x_{n,0}^{k}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
+We pick a resource allocation game (Figure 2) inspired by DSM in electricity grid [64].
+We keep $N=1000$ and $d=24$.
+We choose the action set for each player $n$ as $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$ Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$, and $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$.
+We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$.
 We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{n}^{i}$, i.e., the weighted load at resource $i$.
-The coefficients $a_{n,i}^{(i)}$ are sampled independently from a Gaussian distribution with mean $1$ and variance $0.2$.
-The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are sampled independently from a Gaussian distribution with mean $35$ and variance $5$.
+The coefficients $a_{n,i}^{(i)}$ are i.i.d.
+samples from a Gaussian distribution with mean $1$ and variance $0.2$.
+The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are i.i.d.
+samples from a Gaussian distribution with mean $35$ and variance $5$.
 In Fig.
 2(a), we plot the weighted load for each resource along with their target loads.
-We can observe that the system converges to a Nash equilibrium where the loads are balanced.
+We can observe that the system converges to a NE where the loads are balanced.
 In Fig.
@@
 In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
-We work with $N=100$ players and $d=5$, and action sets $\sum_{i=1}^{d}x_{n}^{i}\leq 10$ for each player $n$.
-To generate a random $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a Gaussian distribution with mean 0 and variance 1.
+We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$.
+Recall that the global cost is given by $\boldsymbol{x}^{\top}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$.
+To generate $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a standard Gaussian distribution.
 We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$.
 A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$.
-The above matrices $\boldsymbol{Q_{n}}$ are obtained in a similar method as $\boldsymbol{G}$ but by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$.
+The above matrices $\boldsymbol{Q_{n}}$ are obtained by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$.
 The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.
@@
 The manager does not observe individual actions, and players do not need to reveal their reward functions, which might even be unknown to them.
-From a technical point of view, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$.
+Technically, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$.
 Our algorithm essentially controls the unknown game such that its NE becomes desirable from an optimization point of view.
@@
 Controlling unknown games with bandit feedback is a new online learning paradigm.
-The game to be controlled can be thought of as a non-stationary but highly structured bandit.
-Viewing game control as a bandit problem can result in both new exciting research questions, and new powerful game control techniques.
-Generalizing our game control setting beyond linear constraints is an exciting new research avenue that can help tackle a wider class of problems.
+The game to be controlled is a non-stationary but highly structured bandit.
+Viewing game control as a bandit problem can result in new exciting research questions, and new powerful game control techniques.
+Generalizing our game control setting beyond linear constraints can help tackle a wider class of problems.

### v3

@@
 Abstract
-Consider a game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager.
-We assume that the game is strongly monotone, so gradient play converges to a unique Nash equilibrium (NE).
-The NE is typically globally inefficient.
-The global performance at NE can be improved by imposing linear constraints on the NE.
-We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE.
-However, this requires knowing the players’ reward functions and action sets.
+Consider a strongly monotone game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager.
+Gradient play converges to a unique Nash equilibrium (NE) that does not optimize the global objective.
+The global performance at NE can be improved by imposing linear constraints on the NE, also known as a generalized Nash equilibrium (GNE).
+We therefore want the manager to control the coefficients such that they impose the desired constraint on the NE.
+However, this requires knowing the players’ rewards and action sets.
 Obtaining this game information is infeasible in a large-scale network and violates user privacy.
@@
 Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets.
-We prove that our algorithm converges with probability 1 to the set of NE that satisfy target linear constraints.
+We prove that our algorithm converges with probability 1 to the set of GNE given by coupled linear constraints.
 We then prove an L2 convergence rate of near-$O(t^{-1/4})$.
 1 Introduction
-Large-scale networks such as in communication, transportation, energy, and computing consist of myriad of agents taking local decisions.
-Agents do not see the bigger picture therefore optimize their local objectives instead.
-The result is often an inefficient equilibrium.
+Large-scale networks such as in communication, transportation, energy, and computing consist of a myriad of agents taking local decisions.
+Agents do not see the bigger picture and instead only optimize their local objectives.
+The result is often an inefficient equilibrium [1].
 Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective.
 However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy.
-As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5].
-Game theory is a natural way to tackle distributed decision-making [6, 7].
-The game structure formalizes the effect of the interaction between users.
+The interaction between the users is formalized as a game [2, 3].
 The players may be cooperative, or they may be selfishly interested in maximizing their reward.
@@
 A Nash Equilibrium (NE) aims to predict the outcome of such a repeated interaction.
-In monotone games [8, 9, 10], which are our focus here, simple distributed algorithms such as gradient descent of players on their reward functions are guaranteed to converge to a NE.
+In monotone games [4, 5, 6], which are our focus here, simple distributed algorithms such as gradient ascent of players on their reward functions are guaranteed to converge to a NE.
 A NE does not optimize global objectives and can lead to poor performance.
@@
 Therefore, in this paper, we focus on the task of imposing linear constraints on the resulting NE of the game the manager controls.
-Linear constraint are common in many optimization formulations.
-When optimizing a global quadratic cost, linear constraint can describe the first-order optimality condition for the NE.
+Linear constraints are common in many optimization formulations.
+When optimizing a global quadratic cost, linear constraints can describe the first-order optimality condition for the NE.
 In resource allocation, linear constraints can describe what should be the ideal loads at each resource at NE.
@@
 In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback.
-Such feedback is easy to monitor in practice and maintains user privacy.
+Such feedback is easy to monitor in practice and preserves user privacy.
 Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
 We prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1.
-We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraint.
-Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping typically assumed to analyze convergence rates [11, 12].
+We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraints.
+Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping that allows for a $O(t^{-1})$ convergence rate [7, 8].
 Apart from stochastic approximation, we employ tools from variational inequalities and the Krasnosel’skiĭ–Mann (KM) algorithm in our analysis.
@@
 1.1 Related Work
-Distributed optimization [13, 14] is concerned with optimizing a target objective distributedly over a network of agents.
-However, these agents are non-interacting computational units so there is no game structure between them.
-To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward function, which is infeasible.
-Moreover, distributed optimization entails significant communication overhead between agents.
-Distributed algorithms that converge to NE have been studied in [15, 16, 17], including in monotone games [18, 19, 20].
+Distributed optimization [9, 10] is concerned with optimizing a target objective over a network of agents.
+However, these agents are non-interacting computational units that are not coupled in a game structure.
+To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward, which is infeasible.
+Moreover, distributed optimization requires significant communication between agents.
+Distributed algorithms that converge to NE have been studied in [11, 12, 13], and in [14, 6, 15] for monotone games.
 The players in our game use simple stochastic gradient play to converge to NE.
-However, if the game is uncontrolled, the NE is typically globally inefficient.
-Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
-Compared to mechanism design approaches [21, 22], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only observes the constraint violation.
+However, if the game is uncontrolled, the NE does not optimize global objectives.
+In our game control scheme, the manager learns to steer the game to an efficient NE, which is a GNE with target linear constraints.
+Convergence to GNE [16] in strongly monotone games is usually studied in the noiseless setting [17, 18, 19], while our learning agents make the convergence to a GNE stochastic.
+Stochastic convergence to GNEs has recently started to attract attention.
+In [20] players share their actions with others, and in [21] the reward function is assumed to be linear.
+Closer to our scenario is [22], which studied a centralized algorithm in a monotone game where the weighted empirical average of the actions is shown to converge in expectation to a GNE.
+In contrast, our GNE is limited to linear coupled constraints in strongly monotone games.
+For this scenario, we provide a distributed stochastic learning scheme that converges in probability 1 to a GNE, without communication between players.
+To the best of our knowledge, our work is the first to study distributed and stochastic convergence to a GNE.
+Our algorithm is a stochastic iterative approach to solving a constrained optimization problem using Lagrange multipliers.
+Convergence guarantees for stochastic augmented Lagrangian schemes are scarce.
+In [23], the continuous-time optimization problem was studied, and [24] provides asymptotic results for a variant of our algorithm.
+We present the convergence rate for our algorithm using novel analysis techniques.
+Unlike mechanism design [25, 26], our players are cooperative or myopic, and our manager does not elicit any players’ private information, but only observes the constraint violation.
 Our approach resembles a Stackelberg game [27, 28, 29], where the manager is the leader and the players are the followers (usually a single follower is considered).
@@
 Our work provides a mechanism that provably shifts the unique NE to a point that satisfies target linear constraints that can force the NE to be efficient (e.g., first-order optimality condition for a global quadratic objective).
-Our learning framework can be viewed as a bandit learning problem where the manager is playing against a bandit that is not stochastic nor fully adversarial [26, 27].
+Our manager is learning with bandit feedback that is not stochastic nor fully adversarial [30, 31].
 Instead, the rewards of this bandit are generated by the game dynamics.
 This “game bandit” introduces a new type of noise.
-The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints.
-Hence, the type of feedback the manager needs concerns the behavior at NE.
+The goal of the manager is to steer the unique NE to a point that satisfies target linear constraints.
+Hence, the feedback the manager needs concerns the behavior at NE.
 However, the dynamics are not at NE every turn.
-In fact, by changing the prices the manager perturbs the system which complicates and delays the convergence to the unique NE.
-Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
-Our problem can be viewed as a Generalized Nash Equilibrium Problem (GNEP), where players seek to converge to a NE subject to coupled constraints [28].
-Efficient computation of GNE in strongly monotone games in the noiseless setting has been widely studied [29, 30, 31].
-Stochastic convergence to GNEs has recently been studied under different assumptions.
-In [32] players share their actions with others, and in [33] the reward function is assumed to be linear.
-In [34], stochastic convergence to GNE is studied more generally, resulting in a centralized algorithm where the weighted empirical average of the actions is shown to converge in expectation.
-We provide a distributed stochastic learning scheme that converges in probability 1 to GNE with linear coupled constraints in strongly monotone games without communication between players.
+In fact, by changing the prices, the manager perturbs the system which complicates and delays the convergence to the unique NE.
+Therefore, constraint violation feedback in this problem suffers from equilibrium noise which is the distance of the dynamics from the time-varying NE.
 Our work belongs to the literature on intervention and control of games [32, 33, 34, 35, 36, 37, 38, 39, 40].
@@
 It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads.
-In [46, 47, 48], a similar approach to ours was considered for discrete congestion games, where the reward function of all players is the same convex function of the number of players that share the chosen resource.
+In [43, 44, 45], a similar approach to ours was considered for discrete congestion games, where the cost function of all players is the same convex function of the number of players that share the chosen resource.
 The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets.
@@
 We consider a more general case where the slower timescale has a non-expansive mapping.
+Our convergence rate is limited by the non-expansive mapping in the slower time-scale, and not by the algorithm players use to update their actions.
+This non-expansive mapping arises due to the projection of the actions into the action sets in the faster time-scale.
+When the action sets are $\mathbb{R}^{d}$, our algorithm becomes a two-time-scale iteration with a contractive mapping in both time-scales, with a convergence rate of $O(t^{-1})$ [7].
+This slowdown is a known challenge in the literature.
+While the optimal bound known for single time-scale fixed point iterations with contractive mappings is $O(t^{-1})$ [8], the best known bound for iterations with non-expansive mappings is $O(t^{-1/3})$ [47].
+Compared to $O(t^{-1/3})$, we attribute the slightly slower convergence rate to the two-time scale dynamics.
 An earlier conference paper [48] was limited to load balancing in resource allocation.
-This paper’s framework is similar to that in [50], but is extended to linear constraint in strongly monotone games.
+This paper extends [48] to linear constraints in strongly monotone games.
 The NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha^{*}}\right)=\boldsymbol{\ell^{*}}$.
-Nevertheless, the results here are completely novel and independent of [50].
-Our convergence guarantees are with probability 1 instead of in L2 like in [50].
+Nevertheless, the results here are novel and independent of [48].
+Our convergence guarantees hold both almost surely and in $L_{2}$, whereas [48] establishes only $L_{2}$ convergence.
 Unlike [48], we provide a finite time convergence error bound.
-Our novel analysis is based on two time-scale stochastic approximation with non-expansive mappings.
+Our novel analysis uses two time-scale stochastic approximation with non-expansive mappings.
 1.2 Outline and Notation
@@
 Section 4 breaks the proofs of our main results into lemmas.
-Section 5 provides numerical simulations for the application described earlier.
+Section 5 provides numerical simulations for the applications described earlier.
 Finally, Section 6 concludes the paper and proposes future research directions.
-[…]
+We use bold letters to denote vectors and matrices, and the standard game-theoretic notation where $\boldsymbol{x}_{-n}$ is the vector of actions for all players except player $n$.
+$\|\cdot\|$ denotes the $\ell_{2}$-norm throughout the paper when applied on a vector, and the operator norm for matrices (under the $\ell_{2}$-norm for vectors).
 2 Problem Formulation
@@
 The utility for each player is defined as:
-$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
+$u_{n}(\boldsymbol{x};\beta_{n}^{1},\ldots,\beta_{n}^{d})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
 Here $r_{n}(\boldsymbol{x})$ is the reward function for player $n$.
-We assume that $r_{n}(\boldsymbol{x})$ is twice continuously differentiable for each $n$ and is greater than or equal to 00 for all $\boldsymbol{x}\in\mathcal{X}$.
+We assume that $r_{n}(\boldsymbol{x})$ is twice continuously differentiable for each $n$.
 The vector $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ denotes the ‘controlled coefficients’ controlled by the game manager.
-We use the same vector notation for $\boldsymbol{\beta}$ and $\boldsymbol{\beta}_{n}$.
+We define the same vector notation for $\boldsymbol{\beta}$ and $\boldsymbol{\beta}_{n}$ as we do for $\boldsymbol{x}$ and $\boldsymbol{x}_{n}$, respectively.
+Such a linear term arises when introducing price per use of a resource, or when agents are cooperative and can compute such a term since they know their own action.
 Figure 1: Control of an unknown game
-The key assumption about the game structure is that it is a strongly monotone game. For this, we define the gradient operator $F(\cdot):\mathbb{R}^{Nd}\mapsto\mathbb{R}^{Nd}$ as
+The key assumption is that the game is strongly monotone. For this, define the gradient operator $F(\cdot):\mathbb{R}^{Nd}\mapsto\mathbb{R}^{Nd}$:
 $F(\boldsymbol{x})\coloneqq(\nabla_{\boldsymbol{x}_{1}}r_{1}(\boldsymbol{x}_{1},\boldsymbol{x}_{-1}),\ldots,\nabla_{\boldsymbol{x}_{N}}r_{N}(\boldsymbol{x}_{N},\boldsymbol{x}_{-N})).$ (2)
 Assumption 1.
-There exists $\mu>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
+There exists $\mu>0$ such that $\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
 $\langle\boldsymbol{y}-\boldsymbol{x},F(\boldsymbol{y})-F(\boldsymbol{x})\rangle\leq-\mu\|\boldsymbol{y}-\boldsymbol{x}\|^{2}.$ (3)
-The above assumption means that the game is a strongly monotone game with parameter $\mu>0$.
+Thus, the game is strongly monotone with parameter $\mu>0$, or equivalently the function $F(\cdot)$ is $\mu$-strongly monotone.
 Monotone games are a well-studied class of games [49].
 The linear term in (1) does not affect the strongly monotone property of the game since the gradient vector becomes $F(\boldsymbol{x})-\boldsymbol{\beta}$.
-Since it is a strongly monotone game, there exists a unique pure NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ corresponding to each vector $\boldsymbol{\beta}$, defined as:
+In this strongly monotone game, there exists a unique pure NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ corresponding to $\boldsymbol{\beta}$, defined as:
 Definition 1.
-A strategy profile $(\boldsymbol{x^{*}}_{n}(\boldsymbol{\beta}),\boldsymbol{x^{*}}_{-n}(\boldsymbol{\beta}))$ is a pure Nash equilibrium (NE) corresponding to coefficients $\boldsymbol{\beta}$, if $u_{n}(\boldsymbol{x}_{n}^{\boldsymbol{*}}(\boldsymbol{\beta}),\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})\geq u_{n}(\boldsymbol{x}_{n},\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})$, for all $\boldsymbol{x}_{n}\in\mathcal{X}$ and all $n\in\mathcal{N}$.
-We make the following additional assumption about the game, which says that the gradient function is Lipschitz continuous.
-It is merely a technical assumption, commonly taken along with the strongly monotone condition [52, 53].
+An action profile $(\boldsymbol{x^{*}}_{n}(\boldsymbol{\beta}),\boldsymbol{x^{*}}_{-n}(\boldsymbol{\beta}))$ is a pure Nash equilibrium (NE) corresponding to coefficients $\boldsymbol{\beta}$, if $u_{n}(\boldsymbol{x}_{n}^{\boldsymbol{*}}(\boldsymbol{\beta}),\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})\geq u_{n}(\boldsymbol{x}_{n},\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})$, for all $\boldsymbol{x}_{n}\in\mathcal{X}$ and all $n\in\mathcal{N}$.
+We assume that the gradients of the players are Lipschitz continuous, which is a standard assumption [50, 51].
 Assumption 2.
-There exists $L>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
+There exists $L>0$ such that $\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
 $\|F(\boldsymbol{y})-F(\boldsymbol{x})\|\leq L\|\boldsymbol{y}-\boldsymbol{x}\|.$ (4)
-Our game evolves in discrete time, so at time $t$ players take actions $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$.
+Our game evolves in discrete time, so at time $t$ player $n$ plays $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$.
 The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the manager’s objective (discussed below (6)).
-Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
-In many scenarios, the players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive (i.e., bandit feedback) or other data they collect.
-The gradient can then be approximated from this data using regression or other supervised machine learning techniques [54].
-These learning schemes make the players’ actions stochastic which adds another type of noise to our analysis.
-We therefore assume that agents observe noisy variants of their gradients, where the noise is a martingale difference sequence with a bounded second moment.
+Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ [4].
+In many scenarios, players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive or other data they collect.
+The gradient can then be approximated from this data [52].
+These learning schemes make the players’ actions stochastic, which adds another type of noise to our analysis.
 Assumption 3.
-Each player $n$ observes $\boldsymbol{g}_{n,t-1}$ at time $t$, where $\boldsymbol{g}_{n,t-1}=\nabla_{\boldsymbol{x}_{n,t-1}}r_{n}(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{n,t}$.
+Each player $n$ observes $\boldsymbol{g}_{n,t-1}$ at time $t$, where $\boldsymbol{g}_{n,t-1}=\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{n,t}$.
 Then $\boldsymbol{g}_{t-1}=F(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{t}$, where $\boldsymbol{M}_{t}$ satisfies
@@
 $\mathcal{F}_{t}=\sigma(\boldsymbol{x}_{\tau},\boldsymbol{\beta}_{\tau}\mid\forall\tau\leq t).$
+The idea behind this assumption is that from the point of view of the players, the rewards are stochastic, i.e., $r_{n}(\boldsymbol{x})\triangleq r_{n}(\boldsymbol{x};\xi)$ for an unknown random variable $\xi$.
+Then the optimal estimator of the gradient of player $n$ from past observations is $\boldsymbol{g}_{n,t}\triangleq\mathbb{E}\left\{\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x}_{t})\mid\mathcal{F}_{t}\right\}$.
+This makes the estimation error $\boldsymbol{M}_{n,t+1}=\boldsymbol{g}_{n,t}-\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x}_{t})$ a Martingale difference sequence, and $\left\|\boldsymbol{M}_{n,t}\right\|$ is bounded if $\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x})$ is bounded.
+Gradient estimation from bandit feedback does not satisfy Assumption 3, which requires unbiased estimates.
 The objective of the manager is to steer the players’ NE towards a point that satisfies a set of $K$ linear constraints, which can together be represented in vector form as
@@
 For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global cost of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$.
-All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$.
+All the action profiles that are optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$.
 Hence, using our approach, we can set $\boldsymbol{A}=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a global optimum.
@@
 In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users.
-Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol.
+Therefore, load balancing is a key control objective for the manager who supervises the resource allocation protocol.
 In the electricity grid, the manager can be the utility company running Demand Side Management (DSM) to cut the production costs of energy at peak hours.
@@
 For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
-In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources.
-We have $d$ linear constraints i.e., $K=d$ and each represents a target load for each resource $i$, i.e.,
 […]
-The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize application-specific objectives such as system efficiency, regulations, and operation costs.
-The controlled coefficients $\boldsymbol{\beta}$ represent incentives to use a resource.
-The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
-A special case is resource allocation without weights, studied in [50].
-Generalizing to a weighted sum of actions allows modeling of power control in wireless networks and load balancing in data centers.
 In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains.
 In data centers, different users send jobs of different sizes to different servers (resources), so the load is a weighted sum of their actions (how many jobs to send where).
+A special case is resource allocation without weights, studied in [48].
 3 Game Control Algorithm
-Algorithm 1 Online Game Control with Bandit Feedback
+Algorithm 1 Online Game Control
 Initialization: Let $\boldsymbol{x}_{0}\in\mathcal{X}$ and $\boldsymbol{\alpha}_{0}\in\mathbb{R}^{K}$.
-Input: Target vector $\boldsymbol{\ell^{*}}$.
+[…]
 For each turn $t\geq 1$ do
   1. 1.
-    The manager observes the vector $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$.
+[…]
   2. 2.
@@
 Depending on the application, this row can represent local parameters known to the player (e.g., the size of the jobs the player sends to the server), or that the manager sends these parameters to players at the beginning of the algorithm.
-The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$.
-To overcome this, we propose an online approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$.
-We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$.
-This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
-Our algorithm is a two time-scale stochastic approximation algorithm.
-In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$.
-In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
-Essentially, our algorithm is an iterative approach to solving the constrained optimization problem using Lagrange multipliers.
-We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
-The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied for all $t$ without communicating, and hence the manager uses an iterative approach instead.
-Similar approaches have been used in [57, 58] for different problems.
-Our players run stochastic gradient descent to optimize their rewards, and therefore are myopic and do not look to manipulate the manager using dynamic strategies.
-This behavior can be encountered with selfish players in a large-scale network or with cooperative players.
-In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [59, 10].
-In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps.
+[…]
+Our algorithm is a two time-scale stochastic approximation iteration.
+In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using gradient play, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$.
+In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action and updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
+Our players run SGD to optimize their rewards, and therefore are myopic and do not manipulate the manager using dynamic strategies.
+This behavior is expected from selfish players in a large-scale network or from cooperative players.
+Selfish players are likely to maximize their rewards using a common algorithm such as gradient ascent [6].
+In particular, when the players do not know their reward functions, small SGD steps allow for gradual learning, as opposed to best-response dynamics that can have large jumps.
 Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
-In a cooperative setting, the behavior of the players is not something to model but to design [60].
+The behavior of cooperative players is not something to model but to design [55].
 The algorithm is used as a distributed protocol programmed into the devices in the network (e.g., WiFi).
@@
 3.1 Performance Guarantees
-The manager wants to converge to a suitable $\boldsymbol{\alpha}$ such that the linear constraints in (6) are satisfied at the NE corresponding to $\boldsymbol{\alpha}$, i.e., converge to the set
+The manager wants to converge to a suitable $\boldsymbol{\alpha}$ such that the linear constraints $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ are satisfied at the NE corresponding to $\boldsymbol{\alpha}$, i.e., converge to the set
 $\mathcal{N}_{opt}=\{\boldsymbol{\alpha}\mid\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}\},$
 where $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ denotes the NE corresponding to $\boldsymbol{\alpha}$.
-We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is just Slater’s condition [61] for convex optimization.
+We assume that Slater’s condition [56] for our problem holds, which allows us to prove that $\mathcal{N}_{opt}$ is non-empty:
 Assumption 4.
-The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
+The set of action profiles is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
 We also make the following standard assumptions on our step-size sequences [57].
 Assumption 5.
-The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ which satisfy:
-  1. (a)
-    $\{\eta_{t}\},\{\epsilon_{t}\}$ are monotonically decreasing sequences that are non-summable, but square summable.
-  2. (b)
-    $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$.
-Our first result shows that $\boldsymbol{\alpha}_{t}$ converges to $\mathcal{N}_{opt}$ and hence $\boldsymbol{A}\boldsymbol{x}_{t}$ converges to $\boldsymbol{\ell^{*}}$.
+The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ and $0.5<\eta,\epsilon<1$ such that $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$.
+The assumption that $0.5<\eta,\epsilon<1$ ensures that the stepsize sequences are monotonically decreasing and square-summable but not summable.
+The assumption that $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$ quantifies the required separation between the time-scales.
+[…]
 Theorem 1.
-Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and
-$\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}},$
-with probability 1.
-The following theorem gives finite-time convergence guarantees for our algorithm.
+Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and with probability 1:
+$\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}.$
+The next theorem gives the convergence rate for our algorithm. The bound holds for all $t>0$ if $T_{1},T_{2}>T_{0}$.
 Theorem 2.
-Under Assumptions 1-5,
-$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right).$
-Therefore the best rate is obtained for $\eta=0.5+\delta/3$ and $\epsilon=0.75+\delta$:
-$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right),$
-where $\delta$ is arbitrarily small.
-The above theorem holds for all $t$ if we choose $T_{1}$ and $T_{2}$ to be sufficiently large and then $\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right)$ for all $t$, for some $C>0$.
-A smaller value of $\epsilon$ gives a better rate.
-The optimal rate is then dictated by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\uparrow\infty}\epsilon_{t}^{2}/\eta_{t}^{3}=0$ ($2\epsilon>3\eta$).
+Under Assumptions 1-5, there exists a $T_{0}>0$ and $C>0$ such that
+$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right),$
+for all $t>T_{0}$.
+Remark 1.
+A smaller value of $\epsilon$ gives a better rate, but it is constrained by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\rightarrow\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$ ($2\epsilon>3\eta$).
+Let $\delta>0$.
+Then for $\eta=0.5+\frac{\delta}{3}$ and $\epsilon=0.75+\delta$ we obtain
+$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right).$
+where $\delta>0$ can be chosen to be arbitrarily small.
 4 Convergence Analysis
@@
 Proof.
-Let $\mathcal{G}(\boldsymbol{\alpha})$ be our $N$-player game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and action set $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$.
-Consider a modified $N$-player game $\mathcal{G}^{\prime}$ with utility function $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ for $n\in\{1,\ldots,N\}$ and action set $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$.
-Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 along with equations (3.3) and (3.6) from [8] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$, and $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
-$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
+Let $\mathcal{G}(\boldsymbol{\alpha})$ be our game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and set of action profiles $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$.
+Consider a modified game $\mathcal{G}^{\prime}$ with utility functions $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ and set of action profiles $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$.
+Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 from [4] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$.
+Further, the Karush-Kuhn-Tucker conditions (equations (3.3) and (3.6) from [4]) state that there exist $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
+$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h_{i}(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
 and
@@
 Let $\boldsymbol{\alpha}=-\boldsymbol{\lambda}$. Then the above equations can be written as
-$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}=0$
+$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h_{i}(\hat{\boldsymbol{x}})\mu_{i}=0$
 and
 $h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\},$
-which implies that $\hat{\boldsymbol{x}}$ is a NE for the game $\mathcal{G}(\boldsymbol{\alpha})$, and hence $\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\hat{\boldsymbol{x}}$.
-We also know that $\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}$.
-Hence there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, showing that $\mathcal{N}_{opt}$ is non-empty.
-One of the main technical challenges in our analysis stems from the constrained actions represented by the convex and compact action sets.
+[…]
+A main technical challenge in our analysis is the convex and compact action sets.
 Indeed, with no constraints on the actions, and under certain assumptions on the matrix $\boldsymbol{A}$, we can show that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is strongly monotone instead of co-coercive (Lemma 4 (b)).
@@
 $g(\boldsymbol{\alpha})=\boldsymbol{\alpha}+\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}}),$
-where $\gamma=\min\{\mu/\|\boldsymbol{A}\|^{2},1\}$ is a constant chosen to satisfy $02\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2}>$ 0.
+where $\gamma=\min\{\mu/\|\boldsymbol{A}\|^{2},1\}$ is a constant chosen to satisfy $2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2}>0$.
 Note that a fixed point of $g(\boldsymbol{\alpha})$ will satisfy our objective of $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$.
@@
 Lemma 4.
-Let $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ be the unique NE given $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. Then,
+Let $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ be the NE given $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. Then,
   1. (a)
@@
 Proof.
-For a fixed $\boldsymbol{\alpha}$, $-F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively.
-Then [51, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
-$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$
-Hence for $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$, we have
+For a fixed $\boldsymbol{\alpha}$, $F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively.
+Then we state two properties of the NE $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$.
+First, [49, Proposition 1.4.2] states that the NE $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ corresponding to $\boldsymbol{\alpha}$ satisfies the variational inequality
+$\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}),\boldsymbol{\alpha}\right)\right\rangle\leq 0$ (11)
+for all $\boldsymbol{x}\in\mathcal{X}$. Second, for a strongly monotone function, [49, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
+$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$ (12)
+Hence for $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$, we have
 $\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|$
-$\leq L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
-=$\stackrel{{\scriptstyle(a)}}{{=}}L^{\prime}\Big{\|}\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
+$\stackrel{{\scriptstyle(b)}}{{=}}L^{\prime}\Big\|\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
 $\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\Big\|$
@@
 $\leq L_{0}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|,$
-where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$. Here equality (a) follows from [51, Proposition 1.5.8] which says
+where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$.
+Here inequality (a) follows from (12) and equality (b) follows from [49, Proposition 1.5.8] which states that the solution of the VI (11) satisfies
 $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})=\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})).$
 This completes the proof for part (a) of Lemma 4.
-Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$ Since $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})$ is the NE corresponding to $\boldsymbol{\alpha}_{1}$, using [51, Proposition 1.4.2], we have for every $\boldsymbol{x}\in\mathcal{X}$ that $\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0$.
-For $\boldsymbol{x}=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$
+Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$. Then for $x=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$ in (11),
 $\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0.$
@@
 $\;\;\;-2\mu\gamma\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
-$\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
+$=\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
 $\stackrel{{\scriptstyle(b)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}.$
@@
 $=\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
-=$\stackrel{{\scriptstyle(a)}}{{=}}\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))$
+$\stackrel{{\scriptstyle(a)}}{{=}}\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))$
 $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1}))\|^{2}$
@@
 $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big\rangle.$ (15c)
-Here equality (a) uses that $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\in\mathcal{X}$ and inequality (b) follows from since the projection operator $\Pi_{\mathcal{X}}(\cdot)$ for any convex set $\mathcal{X}$ is non-expansive.
+Here equality (a) uses that $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\in\mathcal{X}$ and inequality (b) follows since the projection operator $\Pi_{\mathcal{X}}(\cdot)$ for any convex set $\mathcal{X}$ is non-expansive.
 We first simplify the term (15a):
@@
 $\;\;+2\eta_{t}\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]$
-=$\stackrel{{\scriptstyle(a)}}{{=}}\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
+$\stackrel{{\scriptstyle(a)}}{{=}}\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
 $\stackrel{{\scriptstyle(b)}}{{\leq}}\eta_{t}^{2}M_{c}+L_{0}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha}_{t+1}\|^{2}=\eta_{t}^{2}M_{c}+L_{0}^{2}\epsilon_{t}^{2}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|^{2}$
 $\stackrel{{\scriptstyle(c)}}{{\leq}}\eta_{t}^{2}M_{c}+\epsilon_{t}^{2}c_{3}^{2}.$
-Here equality (a) follows from the fact that $\alpha_{t+1}$ is $\mathcal{F}_{t}$ measurable and $\boldsymbol{M}_{t+1}$ is a Martingale difference sequence, hence making the term $\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]=0$.
+Here (a) uses $\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]=0$, since $\alpha_{t+1}$ is $\mathcal{F}_{t}$ measurable and $\boldsymbol{M}_{t+1}$ is a Martingale difference sequence.
 Inequality (b) follows from Lemma 4 part (a), and (c) follows from $\boldsymbol{x}_{t}\in\mathcal{X}$ for some constant $c_{3}$.
@@
 $\stackrel{{\scriptstyle(a)}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|L_{0}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha}_{t+1}\|$
-$\stackrel{{\scriptstyle}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|$
-$\;\;+2\eta_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|$
+[…]
+[…]
 $\stackrel{{\scriptstyle(b)}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}c_{3}+2\eta_{t}\epsilon_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|c_{3}.$
@@
 $\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$
-$\leq(1-\frac{3}{2}\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+\eta_{t}^{2}(2c_{1}^{2}+M_{c})$
+$\leq\left(1-\frac{3}{2}\mu\eta_{t}\right)\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+\eta_{t}^{2}(2c_{1}^{2}+M_{c})$
 $\;\;\;+2\epsilon_{t}^{2}c_{3}^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right].$
@@
 $\;\;\;+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\mathbb{E}[\|p_{t}+q_{t}\|^{2}].$ (18)
-Adding (4.1) and (4.1), we get
+Adding (4.1) and (4.1), and dropping negative terms, we get
 $\mathbb{E}\left[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
@@
 which completes the proof for Lemma 6. ∎
-The above three lemmas are sufficient for proving Theorem 1 (see Appendix).
-We use the properties of two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$.
-This allows us to analyze the slower iteration, i.e., the control updates (8) in the limit $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$.
-We rewrite the iteration as a fixed point iteration with the mapping $g(\cdot)$.
+Lemmas 4-6 are sufficient to prove Theorem 1 (see the Appendix).
+We use two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$.
+Then we can analyze the slower iteration of the control updates (8) in the limit $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$.
+We rewrite (8) as a fixed point iteration with the mapping $g(\cdot)$.
 Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact KM algorithm [56].
@@
 4.2 Finite Time Convergence Guarantees
-In this subsection we detail how to obtain our finite time convergence guarantees.
-This result also uses the non-expansiveness on $g$, but requires a more nuanced argument that bounds the error as follows
+Our finite time convergence guarantees also use the non-expansiveness on $g$, but require a more nuanced argument that bounds the error as follows
 $\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}]$ $\leq 2\mathbb{E}[\|\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))\|^{2}]$
@@
 Lemma 7.
-Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
+Under Assumptions 1-5, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
 $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t}.$
@@
 $+\eta_{t}^{2}(c_{4}+2c_{2}^{2}C_{0}).$
-For $\eta_{t}=(t+T_{1})^{-\eta}$ with $10<\eta<$ 1, this recursion yields [12]
+For $\eta_{t}=(t+T_{1})^{-\eta}$ with $0<\eta<1$, this yields [8]
 $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t},$
 for $t>T_{0}$ for some constant $C^{\prime}$. ∎
-The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$.
-Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$.
-We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) KM algorithm.
-If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [63].
-Since this assumption is not true in our case, we use the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
+[…]
 Lemma 8.
@@
 This inequality follows from Lemma 7 and since $\mathbb{E}[\|q_{t}\|]^{2}\leq\mathbb{E}[\|q_{t}\|^{2}]$.
-Using our assumption that $\epsilon_{t}$ goes to zero, we have that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$.
+Using our assumption that $\epsilon_{t}$ goes to zero, we have that $0<\epsilon_{t}^{\prime}<1/2$ for all $t>T_{0}$.
 For the rest of the proof, we pick $T_{0}$ to be large enough such that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$, $\epsilon_{t}<\gamma/2$ and $\eta_{t}<C_{2}$ for all $t>T_{0}$, for appropriate $C_{1}$ and $C_{2}$.
@@
 $=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}-\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-\|p_{t+1}-p_{t}\|^{2}$
-$\leq-\|p_{t+1}-p_{t}\|^{2}.$
-Here the final inequality follows from the non-expansive nature of $g(\cdot)$. This gives us
+$\leq-\|p_{t+1}-p_{t}\|^{2},$
+where the last inequality uses that $g(\cdot)$ is non-expansive. Then,
 $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\|p_{t+1}-p_{t}\|^{2}-2\langle p_{t+1}-p_{t},q_{t}\rangle$
@@
 $\leq\|p_{t}\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}.$
-Using the assumption that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
+Using the assumption that $0<\epsilon_{t}^{\prime}<1/2$ for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
 $\mathbb{E}[\|p_{i}\|^{2}]$ $\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2\epsilon_{j}^{\prime}\mathbb{E}\|q_{j}\|^{2}$
@@
 The final inequality follows from the fact that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ and hence $2\epsilon\geq 3\eta$.
-Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $14\eta-1>$ 1, so the summation is bounded by a constant.
+Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $4\eta-1>1$, so the summation is bounded by a constant.
 Hence all terms in the right-hand side of (4.2) are bounded by a constant.
@@
 (b) Loads for 3 resources
-Figure 2: Weighted Load Balancing in Resource Allocation Games
+(c) Load balancing for time-varying loads
+Figure 2: Load Balancing in Resource Allocation Games
 (a) Quadratic cost
 (b) Sum of rewards of all players
+(c) Cost when no $\boldsymbol{x}\in\mathcal{X}$ satisfies $2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$
 Figure 3: Quadratic Global Objective
 In this section, we simulate the two applications discussed in Section 2.1.
-In all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$.
-We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player.
+In all experiments, we use the step-size sequences $\eta_{t}=\frac{1}{\left(t+1\right)^{0.501}}$ and $\epsilon_{t}=\frac{1}{\left(t+1\right)^{0.753}}$ since they approximately give the best rate (see Theorem 2).
+To model imperfect gradient information (see Assumption 3), we add Gaussian noise with mean 0 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player.
 We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small.
 In each realization, $\alpha_{0}^{k}$ and $x_{n,0}^{k}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
-We pick a resource allocation game (Figure 2) inspired by DSM in electricity grid [64].
-We keep $N=1000$ and $d=24$.
-We choose the action set for each player $n$ as $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$ Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$, and $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$.
-We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$.
-We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{n}^{i}$, i.e., the weighted load at resource $i$.
-The coefficients $a_{n,i}^{(i)}$ are i.i.d.
-samples from a Gaussian distribution with mean $1$ and variance $0.2$.
-The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are i.i.d.
-samples from a Gaussian distribution with mean $35$ and variance $5$.
+We simulated a demand-side management game (Figure 2) [60] with $N=1000$ and $d=24$ hours.
+The action sets were $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}$, and $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $\hat{\mu}^{i}$ and variance $0.5$.
+Here $\hat{\mu}^{i}$ denote the typical hourly energy consumption limit of a residential customer as given in [60, Fig.
+1].
+Similarly, $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $\hat{\mu}$ and variance $1$, where $\hat{\mu}$ is the typical daily consumption limit.
+Let $s_{i}=\sum_{n=1}^{N}x_{n}^{i}$ be the energy consumption at the $i$-th hour.
+We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $\hat{\omega}^{i}$ and variance $0.5$.
+The value of energy at hour $i$, $\hat{\omega}^{i}$, was adapted from the demand characteristics of residential consumers [61].
+The objectives $s_{i}=\ell_{i}^{*}$ were chosen to shave the peak of the uncontrolled system [60], with the following target loads:
+$\boldsymbol{\ell}^{*}=[$ $53.58,45.09,37.63,34.32,41.16,28.95,37.55,34.18,$
+$27.99,28.15,28.71,28.37,19.96,27.80,23.83,27.93,$
+$20.45,21.44,20.86,28.67,20.89,40.12,37.64,34.32].$
 In Fig.
@@
 2(b), we plot the weighted load for the first 3 resources.
-To study the effectiveness of our algorithm, we also plot the loads in the case of a uncontrolled system., i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$.
+To study the effectiveness of our algorithm, we also plot the loads in the case of an uncontrolled system, i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$.
 It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
 For the quadratic global objective (Figure 3), we consider a quadratic game.
-In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
-We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$.
+In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}^{T}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
+We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid 0\leq\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$.
 Recall that the global cost is given by $\boldsymbol{x}^{\top}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$.
@@
 The above matrices $\boldsymbol{Q_{n}}$ are obtained by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$.
-The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.
+The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way to $\boldsymbol{\rho}$.
+Fig.
+3(a) shows that our algorithm converges to the optimal global cost, which is significantly better than the global cost of the uncontrolled system.
+Even more, Fig.
+3(b) shows an advantage of our scheme over a centralized scheme that optimizes the global cost $\Phi(\boldsymbol{x})$ directly, i.e., where each player updates their action using $\nabla\Phi(\boldsymbol{x})$.
+While both approaches minimize the global cost, our algorithm results in a significantly better sum of rewards $\sum_{n=1}^{N}r_{n}(\boldsymbol{x})$.
+The two plots together show that our algorithm can optimize the global cost under the soft constraint of maintaining a high sum of local player rewards.
+In Figures 2(c) and 3(c), we study the performance of our algorithm beyond our theoretical model and guarantees.
 In Fig.
-3(a), we observe that our algorithm is able to converge to a lower quadratic cost as compared to a system which is left uncontrolled.
+2(c), the target loads $\boldsymbol{\ell^{*}}_{t}$ vary over time.
+The objective now is $\sum_{n=1}^{N}x_{n,t}^{i}=\ell_{i,t}^{*}$.
+Our algorithm easily tracks the time-varying loads under both smoothly varying loads and abruptly changing load profiles.
 In Fig.
-3(b), we plot the sum of rewards for our algorithm.
-This is compared with the sum of rewards if the global objective was directly optimized for, i.e., each player updates their action based on $\nabla\Phi(\boldsymbol{x})$ where $\Phi(\boldsymbol{x})$ is the quadratic cost.
-Our algorithm provides a much higher sum of rewards after convergence.
-The two plots together show that our algorithm can be used as a method to optimize for a global cost while still trying to maximize the rewards for players.
+3(c), we study the case where the linear constraints are infeasible, i.e., $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ is not satisfied by any $x\in\mathcal{X}$.
+In this case, we make a slight modification to our algorithm by projecting the control input $\boldsymbol{\alpha}$ onto a sphere after each iteration.
+This is required to ensure the stability of the algorithm.
+Although it is no longer possible to achieve the optimal cost, our modified algorithm still demonstrates stable behavior and convergence to an action profile with a lower quadratic cost as compared to the uncontrolled system.
 6 Conclusion
-We studied how to control strongly monotone games where the game structure is unknown to the manager.
-The goal of the controller is to ensure the game satisfies target linear constraints at NE.
-Our algorithm only requires the manager to observe the instantaneous constraint violation.
-We prove that our simple online algorithm guarantees convergence to a point $\boldsymbol{x^{*}}$ which satisfies the linear constraints $\boldsymbol{A}\boldsymbol{x^{*}}=\boldsymbol{l}^{*}$ by adjusting control inputs $\boldsymbol{\alpha}$ in real-time.
+We studied how to control a strongly monotone game that is unknown to the manager.
+The goal of the manager is to ensure the game satisfies target linear constraints at NE.
+Our simple algorithm only requires the manager to observe the instantaneous constraint violation.
+We prove that our algorithm guarantees convergence to an $\boldsymbol{x^{*}}$ which satisfies $\boldsymbol{A}\boldsymbol{x^{*}}=\boldsymbol{\ell^{*}}$ by adjusting control inputs $\boldsymbol{\alpha}$ in real-time.
 Further, we show that the convergence rate for our algorithm is $O(t^{-1/4})$.
-Our algorithm is a two time-scale algorithm, where the algorithm does not wait for the players to converge to a NE but uses the feedback from every turn.
+Our two time-scale algorithm does not wait for the players to converge to a NE but uses the feedback from every turn.
 This is an encouraging result since our algorithm is easy to implement and maintains the users’ privacy.
-The manager does not observe individual actions, and players do not need to reveal their reward functions, which might even be unknown to them.
+The manager does not observe individual actions, and players do not reveal their reward functions, which might even be unknown to them.
 Technically, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$.
@@
 The game to be controlled is a non-stationary but highly structured bandit.
-Viewing game control as a bandit problem can result in new exciting research questions, and new powerful game control techniques.
+Viewing game control as a bandit problem can result in new, powerful game control techniques.
 Generalizing our game control setting beyond linear constraints can help tackle a wider class of problems.
+[…]
+7 Proofs from Section 3
+7.1 Proof for Theorem 1
+For a fixed $\boldsymbol{\alpha}$, the ODE limit of iteration (7) is $\dot{\boldsymbol{x}}(t)=F(\boldsymbol{x}(t),\boldsymbol{\alpha})+\mathfrak{b}_{\mathcal{X}}(\boldsymbol{x}(t)).$ Here $\mathfrak{b}_{\mathcal{X}}(\boldsymbol{x}(t))$ is the boundary projection term required to keep $\boldsymbol{x}(t)$ inside $\mathcal{X}$ at all terms.
+Since the game is strongly monotone for any $\boldsymbol{\alpha}$, [4, Theorem 9] shows that for a fixed $\boldsymbol{\alpha}$, there exists a unique globally stable equilibrium of the above ODE, which is the NE $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$.
+Additionally, we have shown that $\|\boldsymbol{\alpha}_{t}\|$ has bounded expectation in Lemma 6 for all $t$, which implies that with probability 1, $\|\boldsymbol{\alpha}_{t}\|<\infty$ for all $t$.
+Then Part (a) of Lemma 4 shows that $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is Lipschitz continuous in $\boldsymbol{\alpha}$.
+Then [58, Lemma 8.1] shows that $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$.
+Now, we can write iteration (8) as
+$\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t})+\epsilon_{t}^{\prime}\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})),$
+where $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$. The ODE limit for this iteration is
+$\dot{\boldsymbol{\alpha}}(t)=g(\boldsymbol{\alpha}(t))-\boldsymbol{\alpha}(t).$ (22)
+The set of fixed points of $g(\cdot)$ is $\mathcal{N}_{opt}$, which is non-empty as shown in Lemma 3.
+Let $\boldsymbol{\alpha^{*}}$ be any such point.
+[62, Theorem 3.1] shows that the solutions of $\dot{\boldsymbol{y}}(t)=f(\boldsymbol{y}(t))-\boldsymbol{y}(t)$ converge to the set of fixed points of $f(\cdot)$, when $f$ is non-expansive under $\ell_{2}$ norm and the set of fixed points of $f(\cdot)$ is non-empty.
+Since $\boldsymbol{\alpha}(t)$ converges to a point in $\mathcal{N}_{opt}$, this completes our proof that the iterates $\boldsymbol{\alpha}_{t}$ converge to $\mathcal{N}_{opt}$.
+7.2 Proof for Theorem 2
+Let $T_{0}$ be the time after which the bounds in Lemma 7 and Lemma 8 hold. Then for $t>T_{0}$,
+$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}]$
+$\leq 2\mathbb{E}[\|\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))\|^{2}]+2\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]$
+$\leq 2\|\boldsymbol{A}\|^{2}\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+2\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]$
+$\leq C\left(\eta_{t}+1/(t\epsilon_{t})\right),$
+for $C=\max\{2\|\boldsymbol{A}\|^{2}C^{\prime},2C^{\prime\prime}\}$.